### v2

# Certain conjectures of Clozel and Shahidi on Arthur packets

Abstract.
In this paper, applying the intersection theory of local Arthur packets, for symplectic and split odd special orthogonal groups $\mathrm{G}_{n}$, we prove the enhanced Shahidi conjecture on generic representations in local Arthur packets. As an application, we show that given any automorphic representation in the discrete spectrum of $\mathrm{G}_{n}(\mathbb{A})$, if a finite local component is generic, then almost all local components are generic. Next, we give a classification of the unramified representations of Arthur type for $\mathrm{G}_{n}$ and show that any unramified representation of $\mathrm{G}_{n}$ of Arthur type lies in exactly one local Arthur packet, moreover, it lies in the $L$-packet associated to an anti-generic local Arthur packet. As applications, we prove a reciprocal conjecture of Clozel and confirm a conjecture of Arthur and Clozel on the local components of automorphic representations in the discrete spectrum of $\mathrm{G}_{n}$.
Key words and phrases:
Endoscopic Classification, Local Arthur Packets, Local Arthur Parameters, Intersection of Local Arthur Packets, Enhanced Shahidi Conjecture, Clozel Conjecture
2020 Mathematics Subject Classification:
Primary 11F70, 22E50; Secondary 11F85, 22E55
The research of the second-named author has been partially supported by the NSF Grants DMS-1702218, DMS-1848058
1. Introduction
Let $k$ be a number field and $\mathbb{A}_{k}$ is the ring of adeles. Let $\mathrm{G}$ be a connected reductive group defined over $k$. A main theme in the theory of automorphic forms is to study the discrete spectrum of $\mathrm{G}(\mathbb{A}_{k})$. In [Art89, §6, §8], Arthur conjectured that there are local and global Arthur packets theories for $\mathrm{G}$, parameterized by local and global Arthur parameters, such that the discrete spectrum of $\mathrm{G}(\mathbb{A}_{k})$ is a union of global Arthur packets. More precisely, let $L_{k}$ be the hypothetical Langlands group, then the set of global Arthur parameters $\Psi(\mathrm{G}(k))$ consists of admissible homomorphisms as follows:
$\psi:L_{k}\times\mathrm{SL}_{2}^{A}(\mathbb{C})\rightarrow{}^{L}\mathrm{G}(k),$
where ${}^{L}\mathrm{G}(k)=\widehat{\mathrm{G}}(\mathbb{C})\rtimes W_{k}$ is the global Langlands $L$-group of $\mathrm{G}$, $\widehat{\mathrm{G}}(\mathbb{C})$ is the Langlands dual group of $\mathrm{G}$ and $W_{k}$ is the Weil group of $k$. Let $\Pi_{\psi}$ be the global Arthur packet corresponding to a global Arthur parameter $\psi$.
For any local place $v$, the set of local Arthur parameters $\Psi^{+}(\mathrm{G}(k_{v}))$ consists of admissible homomorphisms as follows:
$\psi_{v}:L_{k_{v}}\times\mathrm{SL}_{2}^{A}(\mathbb{C})\rightarrow{}^{L}\mathrm{G}(k_{v}),$
where ${}^{L}\mathrm{G}(k_{v})=\widehat{\mathrm{G}}(\mathbb{C})\rtimes W_{k_{v}}$ is the local Langlands $L$-group of $\mathrm{G}$, $W_{k_{v}}$ is the Weil group of $k_{v}$, and
$L_{k_{v}}=\begin{cases}W_{k_{v}},&v\text{ Archimedean,}\\ W_{k_{v}}\times\mathrm{SL}_{2}^{D}(\mathbb{C}),&v\text{ non-Archimedean.}\end{cases}$
Here, $\psi_{v}|_{W_{k_{v}}}=\bigoplus_{i}\phi_{i}|\cdot|^{x_{i}}$, where $\phi_{i}(W_{k_{v}})$ is bounded and consists of semi-simple elements, $x_{i}\in\mathbb{R}$ and $|x_{i}|<\frac{1}{2}$. For each local Arthur parameter $\psi_{v}$, Arthur associated a local $L$-parameter $\phi_{\psi_{v}}$ as follows
(1.1) $\phi_{\psi_{v}}(w,x)=\psi_{v}\left(w,x,\begin{pmatrix}|w|^{\frac{1}{2}}&0\\ 0&|w|^{-\frac{1}{2}}\\ \end{pmatrix}\right).$
A local Arthur parameter $\psi_{v}$ is called generic if $\psi_{v}|_{\mathrm{SL}_{2}^{A}(\mathbb{C})}$ is trivial; and is called tempered if additionally $x_{i}=0$ for all $i$. Let $\Psi(\mathrm{G}(k_{v}))$ be the subset of local Arthur parameters with $x_{i}=0$ for all $i$. Let $\Pi_{\psi_{v}}$ be the local Arthur packet corresponding to a global Arthur parameter $\psi_{v}$. By Arthur’s conjectures, given any $\pi\in\Pi_{\psi}$, we have $\pi=\otimes_{v}\pi_{v}$ and for each local place $v$, $\pi_{v}\in\Pi_{\psi_{v}}$. Here $\psi_{v}$ is the restriction of $\psi$ to $L_{k_{v}}$. In [Art13], Arthur proved these local and global conjectures for symplectic and quasi-split special orthogonal groups.
There are many conjectures related to the local Arthur packets $\Pi_{\psi_{v}}$. In this paper, applying the results in [HLL22] on the intersection theory of local Arthur packets, we focus on
  1. (1)
    the enhanced Shahidi conjecture on generic representations in local Arthur packets for quasi-split connected reductive groups; and,
  2. (2)
    the Clozel conjectures on unramified representations in local Arthur packets and on local components of automorphic representations in the discrete spectrum.
1.1. The enhanced Shahidi conjecture
Let $\mathrm{G}$ be a quasi-split connected reductive group defined over a non-Archimedean local field $k_{v}$ and let $G=\mathrm{G}(k_{v})$. The well-known Shahidi conjecture states that tempered local $L$-packets of $G$ have generic members ([Sha90, Conjecture 9.4]). This conjecture has been generalized by Jiang ([Jia14], see also [LS22, Conjecture 1.6]) to non-tempered local Arthur packets. Recently, Shahidi made an enhanced conjecture as follows.
Conjecture 1.1 (Conjecture 4.4, [LS22, Conjecture 1.5], Enhanced Shahidi Conjecture).
For any quasi-split connected reductive group $G$, assume that there is a local Arthur packets theory for $G$ as conjectured in [Art89, §6].
  1. (1)
    For any local Arthur parameter $\psi\in\Psi(G)$, the local Arthur packet $\Pi_{\psi}$ is tempered if and only if it has a generic member.
  2. (2)
    For any local Arthur parameter $\psi\in\Psi^{+}(G)$, the local Arthur packet $\Pi_{\psi}$ is generic if and only if it has a generic member.
In [LS22], the second-named author and Shahidi proved Conjecture 1.1 for symplectic and quasi-split special orthogonal groups, applying the matching method of endoscopic liftings, assuming properties of the wavefront sets of certain bitorsor representations.
Let $\mathrm{G}_{n}$ be the symplectic or the split odd special orthogonal group of rank $n$ and let $G_{n}=\mathrm{G}_{n}(k_{v})$. As an application of the authors’ results on the intersection theory of local Arthur packets in [HLL22] and the classification of generic representations as in [Mui98, JS04, Liu11], we prove Conjecture 1.1 for $G_{n}$, without any assumption.
Theorem 1.2 (Theorem 4.7).
Conjecture 1.1 is valid for $G_{n}$. In particular, any irreducible tempered generic representation of $G_{n}$ lies in exactly one local Arthur packet.
We remark that the method used here is different from that in [LS22]. Later, in [HLLZ22], jointly with Zhang, we gave another proof of the enhanced Shahidi conjecture for symplectic and split odd special orthogonal groups, applying the closure ordering relations among the local $L$-parameters of representations in local Arthur packets, which is geometric in nature. In particular, this provides a framework of proving the enhanced Shahidi conjecture for general quasi-split connected reductive groups.
As an application of Theorem 1.2, we obtain the following result on local components of automorphic representations in global Arthur packets with a generic local component at some finite place. This result has its own interest and plays a role towards the generalized Ramanujan conjecture of $\mathrm{G}_{n}$ (see Theorem 1.5 below).
Theorem 1.3 (Theorem 6.2).
Suppose that $\pi=\otimes\pi_{v}$ lies in a global Arthur packet $\Pi_{\psi}$ of $\mathrm{G}_{n}(\mathbb{A}_{k})$ and that there exists a finite place $v_{0}$ such that $\pi_{v_{0}}$ is generic. Then the followings hold:
  1. (1)
    The global Arthur parameter is of the form $\psi=\boxplus_{i=1}^{m}\mu_{i}\otimes S_{1},$ where $\mu_{i}$’s are irreducible cuspidal representations of general linear groups.
  2. (2)
    For any place $v,$ $\psi_{v}$ is generic. In particular, $\pi_{v}$ is generic for almost all finite places $v$.
  3. (3)
    Furthermore, for almost all places $v$, the local Arthur parameter $\psi_{v}$ has trivial restrictions to both the Deligne-$\mathrm{SL}_{2}(\mathbb{C})$ and Arthur-$\mathrm{SL}_{2}(\mathbb{C}).$
Based on Theorem 1.3, we propose the following conjecture for general connected reductive groups.
Conjecture 1.4 (Conjecture 6.12).
Let $\pi$ be an automorphic representation in the discrete spectrum of a connected reductive group $\mathrm{G}(\mathbb{A}_{k})$. Suppose there exists a finite place $v_{0}$ of $k$ such that $\pi_{v_{0}}$ is generic, then $\pi_{v}$ is generic for almost all places.
Note that when $\mathrm{G}=\mathrm{G}_{n}$, Conjecture 1.4 is exactly Part (2) of Theorem 1.3. We remark that in Conjecture 1.4 and Theorem 1.3(2), we do not assume the Ramanujan conjecture of $\mathrm{GL}_{n}$. On the other hand, if we do assume the Ramanujan conjecture of $\mathrm{GL}_{n}$, then we obtain a stronger conclusion as follows.
Theorem 1.5 (Theorem 6.13).
Assume the Ramanujan conjecture for $\mathrm{GL}_{n}$. Suppose that $\pi=\otimes\pi_{v}$ lies in a global Arthur packet $\Pi_{\psi}$ of $\mathrm{G}_{n}({\mathbb{A}}_{k})$ and that there exists a finite place $v_{0}$ such that $\pi_{v_{0}}$ is generic. Then $\pi_{v}$ is tempered for all places $v$.
This is towards the generalized Ramanujan problem, as posted by Sarnak in [Sar05], to characterize the failure of temperedness of local components of general automorphic representations. Note that, applying the Langlands functoriality, it is known that the Ramanujan conjecture for $\mathrm{GL}_{n}$ implies the generalized Ramanujan conjecture for quasi-split classical groups $\mathrm{G}$, that is, if $\pi=\otimes_{v}\pi_{v}$ is a globally generic cuspidal automorphic representation of $\mathrm{G}({\mathbb{A}}_{k})$, then $\pi_{v}$ are all tempered (see [Sha11, (6.6)]). Theorem 1.5 improves the results in [Sha11, Theorem 6.1 and Conjecture 6.5], where locally generic (at every local place) condition is assumed.
We remark that Clozel made the following conjecture.
Conjecture 1.6 ([Clo07, Conjecture 5]).
Let $\pi$ be an automorphic representation in the discrete spectrum of a reductive group $\mathrm{G}({\mathbb{A}}_{k})$. Suppose there exists a finite place $v_{0}$ of $k$ such that $\pi_{v_{0}}$ is a Steinberg representation, then $\pi_{v}$ is tempered for all places $v$.
Conjecture 1.6 is also true under the assumption of the Ramanujan conjecture for $\mathrm{GL}_{n}$, as a direct corollary of Theorem 1.5, since Steinberg representations are generic.
1.2. Unramified representations of Arthur type and the Clozel conjectures
Considering the importance of unramified representations in the theory of automorphic forms and automorphic representations, it is desirable to study more closely the unramified representations of Arthur type. In general, we expect a dual version of the enhanced Shahidi conjecture (Conjecture 1.1) as follows.
Conjecture 1.7 (Conjecture 5.1).
Let $\mathrm{G}$ be a connected reductive group defined over a non-Archimedean local field $k_{v}$ and let $G=\mathrm{G}(k_{v})$. Assume that there is a local Arthur packets theory for $G$ as conjectured in [Art89, §6]. Then any unramified representation of $G$ of Arthur type lies in exactly one local Arthur packet. Moreover, it lies in the $L$-packet associated to an anti-generic local Arthur packet.
Again, let $\mathrm{G}_{n}$ be the symplectic or the split odd special orthogonal group of rank $n$ and let $G_{n}=\mathrm{G}_{n}(k_{v})$. In §5, we give a characterization of unramified representations of $G_{n}$ of Arthur type in terms of their $L$-data (see Theorem 5.2), making use of Algorithm 3.9. As an application, we prove Conjecture 1.7 for $G_{n}$ (see Corollary 5.3). We remark that Muić and Tadić ([MT11, Theorem 0-5]) showed that any negative unramified representation is of Arthur type and lies in the $L$-packet associated to an anti-tempered local Arthur packet. Theorem 5.2 shows that any unramified representation of Arthur type is negative, based on which, Conjecture 1.7 can also be implied by [MT11, Theorem 0-5].
Conjecture 1.7 is the reciprocal of the following conjecture of Clozel.
Conjecture 1.8 ([Clo07, Conjecture 2A]).
Let $\mathrm{G}$ be a connected reductive group defined over a non-Archimedean local field $k_{v}$ and let $G=\mathrm{G}(k_{v})$. Assume that there is a local Arthur packets theory for $G$ as conjectured in [Art89, §6]. Suppose that $\psi$ is a local Arthur parameter of $G$ such that $\phi_{\psi}$ is unramified. Then the local Arthur packet $\Pi_{\psi}$ contains a unique unramified representation. More specifically, the unramified representation is the one associated to $\phi_{\psi}$ via the Satake isomorphism (((([Sat63])))).
We remark that the first part of Conjecture 1.8 has already been proved by Mœglin ([Mœ09b, Proposition 6.4]) for classical groups. As an application of the results in §5, we fully prove Conjecture 1.8 (see Theorem 6.9). We remark that although it is not stated explicitly, one may conclude the second part of Conjecture 1.8 from [Mœ09b] as well, using different methods from this paper. As a corollary, we confirm a conjecture of Arthur and Clozel as follows (see [Sha11, Conjecture 6.1]).
Theorem 1.9.
Let $\pi=\otimes_{v}\pi_{v}$ be an automorphic representation of $\mathrm{G}_{n}(\mathbb{A}_{k})$ in the discrete spectrum with global Arthur parameter $\psi\in\Psi(\mathrm{G}_{n}(k))$. Then, for almost all finite places $v$, we have $\pi_{v}\in\Pi_{\phi_{\psi_{v}}}.$
Clozel also made the following deep conjectures on local components of automorphic representations in the discrete spectrum of connected reductive groups $\mathrm{G}$, towards the generalized Ramanujan problem as posted by Sarnak.
Conjecture 1.10 ([Clo07, Conjectures 2, 4]).
Let $\mathrm{G}$ be a connected reductive group defined over a number field $k$. Assume that there are local and global Arthur packets theories for $\mathrm{G}$ as conjectured in [Art89, §6, §8]. Let $\pi=\otimes_{v}\pi_{v}$ be an automorphic representation of $\mathrm{G}(\mathbb{A}_{k})$ in the discrete spectrum with global Arthur parameter $\psi\in\Psi(\mathrm{G}(k))$.
  1. (1)
    Then, for any finite place $v$ such that $\pi_{v}$ is unramified, the Satake parameter of $\pi_{v}$ is of the form $\phi_{\psi_{v}}(\mathrm{Frob}_{v})$ for some $\psi_{v}\in\Psi(\mathrm{G}(k_{v}))$.
  2. (2)
    If there exists a finite place ${v_{0}}$ such that $\pi_{v_{0}}$ is unramified and tempered, then every component of $\pi$ is tempered.
  3. (3)
    In general, for any finite place $v$ such that $\pi_{v}$ is unramified, write the multi-set of the absolute value of eigenvalues of $\phi_{\psi_{v}}(\mathrm{Frob}_{v})$ as $\{q_{v}^{w_{v,1}},\ldots,q_{v}^{w_{v,N}}\}$, where $q_{v}$ is the cardinality of the residue field of $k_{v}$. Then the multi-set $\{w_{v,1},\ldots,w_{v,N}\}$ is independent of the unramified place $v$.
We remark that the key point of Conjecture 1.10(1) is that in the statement, $\psi_{v}\in\Psi(\mathrm{G}(k_{v}))$, not $\Psi^{+}(\mathrm{G}(k_{v}))$. As applications of the results in §5, we show that the Ramanujan conjecture of $\mathrm{GL}_{n}$ implies Conjecture 1.10 for the groups $\mathrm{G}_{n}$ (see Theorems 6.7 and 6.11). This illustrates the depth and difficulties of these conjectures. Clozel originally made these Conjectures for $L^{2}$-automorphic representations, here for simplicity, we only considered the case of discrete spectrum. Note that by Langlands’ theory of Eisenstein series, the corresponding conjectures for $L^{2}$-automorphic representations reduce to those for the discrete spectrum.
Following is the structure of this paper. In §2, we recall necessary notation and preliminaries. In §3, we recall Atobe’s reformulation on Mœglin’s construction of local Arthur packets, and certain algorithms and results from the intersection theory of local Arthur packets. In §4, we study the existence of generic representations in local Arthur packets and prove Conjecture 1.1 for $\mathrm{G}_{n}$. In §5, we classify unramified representations of $\mathrm{G}_{n}$ of Arthur type in terms of their $L$-data, and we verify Conjecture 1.7 for $\mathrm{G}_{n}$. In §6, we discuss the global implications of the results in §4, 5.
Acknowledgements
The authors would like to thank Dihua Jiang and Freydoon Shahidi for their interests, constant support, helpful discussions, and for bringing our attention on the topics of unramified representations in §5, 6. The authors also would like to thank Hiraku Atobe for helpful communications, comments, and suggestions, and for generously sharing his sage code on the reformulation of Mœglin construction of local Arthur packets. Finally, the authors would like to thank Wee Teck Gan, Tasho Kaletha, Anantharam Raghuram, David Renard, Bin Xu, and Lei Zhang for helpful comments and suggestions.
2. Notation and preliminaries
Let $F$ be a local non-Archimedean field of characteristic 00 with normalized absolute value ||.$\cdot$|.| ⋅ | . We also regard ||$\cdot$|| ⋅ | as a character of $\mathrm{GL}_{n}(F)$ via composition with the determinant. Set $G_{n}$ to be split $\mathrm{SO}_{2n+1}(F)$ or $\mathrm{Sp}_{2n}(F).$ We write $\Pi(G)$ for the set of equivalence classes of irreducible smooth representations of a group $G.$ We assume that every representation is smooth.
Suppose $\Pi_{1},\Pi_{2}$ are representations of finite length. We denote $[\Pi_{1}]$ the image of $\Pi_{1}$ in the Grothendieck group. We write $\Pi_{1}\geq\Pi_{2}$ if $[\Pi_{1}]-[\Pi_{2}]$ is a non-negative linear combination of irreducible representations.
For a multi-set $X$ and $a\in X$, we let $m_{X}(a)$ denote the multiplicity of $a$ in $X.$ Let $Y$ be another multi-set. We define the following multi-sets $Z$ by specifying the multiplicity $m_{Z}(a)$ for each $a$ such that $m_{X}(a)+m_{Y}(a)>0$.
  1. $\bullet$
    The sum of multi-sets $X+Y$: $m_{X+Y}(a):=m_{X}(a)+m_{Y}(a)$.
  2. $\bullet$
    The union of multi-sets $X\cup Y$: $m_{X\cup Y}(a):=\max(m_{X}(a),m_{Y}(a)).$
  3. $\bullet$
    The difference of multi-sets $X\setminus Y$: $m_{X\setminus Y}(a):=\max(m_{X}(a)-m_{Y}(a),0).$
  4. $\bullet$
    The intersection of multi-sets $X\cap Y$: $m_{X\cap Y}(a):=\min(m_{X}(a),m_{Y}(a)).$
  5. $\bullet$
    The symmetric difference of multi-sets $X\Delta Y:=(X\cup Y)\setminus(X\cap Y).$
2.1. Langlands classification
In this subsection, we recall the Langlands classification for the groups $\mathrm{GL}_{n}(F)$ and $G_{n}$ (see [Kon03] for a general setting). Let $n$ be a positive integer and fix a Borel subgroup of $\mathrm{GL}_{n}(F)$. Let $P$ be a standard parabolic subgroup of $\mathrm{GL}_{n}(F)$ with Levi subgroup $M\cong\mathrm{GL}_{n_{1}}(F)\times\cdots\times\mathrm{GL}_{n_{r}}(F).$ Let $\tau_{i}\in\Pi(\mathrm{GL}_{n_{i}}(F))$ for $i=1,2,\dots,r.$ We set
$\tau_{1}\times\cdots\times\tau_{r}:=\mathrm{Ind}_{P}^{\mathrm{GL}_{n}(F)}(\tau_{1}\otimes\cdots\otimes\tau_{r})$
to be the normalized parabolic induction. We define a segment $[x,y]_{\rho}$ to be a set of supercuspidal representations of the form
$[x,y]_{\rho}=\{\rho|\cdot|^{x},\rho|\cdot|^{x-1},\dots,\rho|\cdot|^{y}\},$
where $\rho$ is an irreducible unitary supercuspidal representation of $\mathrm{GL}_{n}(F)$ and $x,y\in\mathbb{R}$ such that $x-y$ is a non-negative integer. We denote a Steinberg representation attached to the segment $[x,y]_{\rho}$ by $\Delta_{\rho}[x,y].$ This is the unique irreducible subrepresentation of $\rho|\cdot|^{x}\times\cdots\times\rho|\cdot|^{y}.$ It is an essentially discrete series representation of $\mathrm{GL}_{n(x-y+1)}(F).$ When it is clear in context, we refer to both $[x,y]_{\rho}$ and $\Delta_{\rho}[x,y]$ as segments. We also set $Z_{\rho}[y,x]$ to be the unique irreducible quotient of $\rho|\cdot|^{x}\times\cdots\times\rho|\cdot|^{y}.$ In the case $y=x+1,$ we set $\Delta_{\rho}[x,x+1]=Z_{\rho}[x+1,x]$ to be the trivial representation of $\mathrm{GL}_{0}(F).$
The Langlands classification for $\mathrm{GL}_{n}(F)$ states that any irreducible representation $\tau$ of $\mathrm{GL}_{n}(F)$ can be realized as a unique irreducible subrepresentation of a parabolic induction of the form
$\Delta_{\rho_{1}}[x_{1},y_{1}]\times\cdots\times\Delta_{\rho_{r}}[x_{r},y_{r}],$
where $\rho_{i}$ is an irreducible unitary supercuspidal representation of $\mathrm{GL}_{n_{i}}(F),$ $[x_{i},y_{i}]_{\rho_{i}}$ is a segment, and $x_{1}+y_{1}\leq\cdots\leq x_{r}+y_{r}.$ In this setting, we write
$\tau=L(\Delta_{\rho_{1}}[x_{1},y_{1}],\dots,\Delta_{\rho_{r}}[x_{r},y_{r}]).$
Let $(x_{i,j})_{1\leq i\leq s,1\leq j\leq t}$ be real numbers such that $x_{i,j}=x_{1,1}-i+j.$ We define a (shifted) Speh representation to be the irreducible representation given by
$\begin{pmatrix}x_{1,1}&\cdots&x_{1,t}\\ \vdots&\ddots&\vdots\\ x_{s,1}&\cdots&x_{s,t}\end{pmatrix}_{\rho}:=L(\Delta_{\rho}[x_{1,1},x_{s,1}],\dots,\Delta_{\rho}[x_{1,t},x_{s,t}]).$
Fix a Borel subgroup of $G_{n}$ and let $P$ be a standard parabolic subgroup of $G_{n}$ with Levi subgroup $M\cong\mathrm{GL}_{n_{1}}(F)\times\cdots\times\mathrm{GL}_{n_{r}}(F)\times G_{m}.$ Let $\tau_{i}$ be a representation of $\mathrm{GL}_{n_{i}}(F)$ for $i=1,2,\dots,r$ and $\sigma$ be a representation of $G_{m}$. We set
$\tau_{1}\times\cdots\times\tau_{r}\rtimes\sigma:=\mathrm{Ind}_{P}^{G_{n}}(\tau_{1}\otimes\cdots\otimes\tau_{r}\otimes\sigma)$
to be the normalized parabolic induction.
The Langlands classification for $G_{n}$ states that every irreducible representation $\pi$ of $G_{n}$ is a unique irreducible subrepresentation of
$\Delta_{\rho_{1}}[x_{1},y_{1}]\times\cdots\times\Delta_{\rho_{r}}[x_{r},y_{r}]\rtimes\pi_{temp},$
where $\rho_{i}$ is an irreducible unitary supercuspidal representation of $\mathrm{GL}_{n_{i}}(F),$ $x_{1}+y_{1}\leq\cdots\leq x_{r}+y_{r}<0,$ and $\pi_{temp}$ is an irreducible tempered representation of $G_{m}.$ In this case, we write
$\pi=L(\Delta_{\rho_{1}}[x_{1},y_{1}],\dots,\Delta_{\rho_{r}}[x_{r},y_{r}];\pi_{temp}),$
and call $(\Delta_{\rho_{1}}[x_{1},y_{1}],\dots,\Delta_{\rho_{r}}[x_{r},y_{r}];\pi_{temp})$ the Langlands data, or $L$-data, of $\pi.$ In §2.4, We give more detailed parametrization of the tempered representation $\pi_{temp}$ using Arthur’s theory.
2.2. Derivatives and socles
Let $\pi$ be a smooth representation of $G_{n}$ of finite length. We let $Jac_{P}(\pi)$ be the Jacquet module of $\pi$ with respect to a parabolic subgroup $P$ of $G_{n}.$ We denote the semisimplification of $Jac_{P}(\pi)$ by $[Jac_{P}(\pi)].$
Definition 2.1.
Let $P_{d}$ be a standard parabolic subgroup of $G_{n}$ with Levi subgroup isomorphic to $\mathrm{GL}_{d}(F)\times G_{n-d},$ $x\in\mathbb{R},$ and $\rho$ be an irreducible unitary self-dual supercuspidal representation of $\mathrm{GL}_{d}(F).$ We define the $\rho|\cdot|^{x}$-derivative of $\pi$, denoted $D_{\rho|\cdot|^{x}}(\pi),$ to be a semisimple representation satisfying
$[Jac_{P_{d}}(\pi)]=\rho|\cdot|^{x}\otimes D_{\rho|\cdot|^{x}}(\pi)+\sum_{i}\tau_{i}\otimes\pi_{i},$
where the sum is over all irreducible representations $\tau_{i}$ of $\mathrm{GL}_{d}(F)$ such that $\tau_{i}\not\cong\rho|\cdot|^{x}.$
We set $D_{\rho|\cdot|^{x}}^{(0)}(\pi)=\pi$ and for $k$ a non-negative integer, we define recursively
$D_{\rho|\cdot|^{x}}^{(k)}(\pi)=\frac{1}{k}D_{\rho|\cdot|^{x}}\circ D_{\rho|\cdot|^{x}}^{(k-1)}(\pi).$
If $D_{\rho|\cdot|^{x}}^{(k)}(\pi)\neq 0$, but $D_{\rho|\cdot|^{x}}^{(k+1)}(\pi)=0,$ then we say that $D_{\rho|\cdot|^{x}}^{(k)}(\pi)$ is the highest $\rho|\cdot|^{x}$-derivative of $\pi$. If $D_{\rho|\cdot|^{x}}(\pi)=0,$ then say that $\pi$ is $\rho|\cdot|^{x}$-reduced.
We also need to define derivatives for $\mathrm{GL}_{n}(F).$ However, in this situation, we must distinguish between left and right derivatives. We follow [Xu17a, §5].
Definition 2.2.
Let $P_{d}$ (resp. $Q_{d}$) be a standard parabolic subgroup of $\mathrm{GL}_{n}(F)$ with Levi subgroup isomorphic to $\mathrm{GL}_{d}(F)\times\mathrm{GL}_{n-d}(F)$ (resp. $\mathrm{GL}_{n-d}(F)\times\mathrm{GL}_{d}(F)$), $x\in\mathbb{R},$ $\sigma$ be a smooth representation of $\mathrm{GL}_{n}(F)$, and $\rho$ be an irreducible unitary self-dual supercuspidal representation of $\mathrm{GL}_{d}(F).$ We define the left (resp. right) $\rho|\cdot|^{x}$-derivative of $\sigma$, denoted $D_{\rho|\cdot|^{x}}(\sigma)$ (resp. $D_{\rho|\cdot|^{x}}^{op}(\sigma)$), to be a semisimple representation satisfying
$[Jac_{P_{d}}(\sigma)]=\rho|\cdot|^{x}\otimes D_{\rho|\cdot|^{x}}(\sigma)+\sum_{i}\tau_{i}\otimes\sigma_{i},$
[…]
where the sum is over all irreducible representations $\tau_{i}$ of $\mathrm{GL}_{d}(F)$ such that $\tau_{i}\not\cong\rho|\cdot|^{x}.$
Note that the right derivative defined in [Xu17a, §5] uses the contragredient; however, for our purposes, we are only concerned with derivatives for self-dual $\rho.$ Atobe and Mínguez gave explicit formulas for computing the highest nonzero derivative in the good parity case ([AM23, Proposition 6.1, Theorem 7.1]). We note that these derivatives satisfy the Leibniz rules below.
Lemma 2.3 ([Xu17a, §5]).
Let $\rho$ be an irreducible unitary self-dual supercuspidal representation of $\mathrm{GL}_{d}(F)$ and $x\in\mathbb{R}.$
  1. 1.
    For $\sigma,\sigma_{1},\sigma_{2}\in\Pi(\mathrm{GL}_{n}(F)),\ \tau\in\Pi(G_{m})$, we have
    $D_{\rho|\cdot|^{x}}(\sigma\rtimes\tau)=D_{\rho|\cdot|^{x}}(\sigma)\times\tau+D_{\rho|\cdot|^{-x}}^{op}(\sigma)\times\tau+\sigma\rtimes D_{\rho|\cdot|^{x}}(\tau),$
    $D_{\rho|\cdot|^{x}}(\sigma_{1}\times\sigma_{2})$ $=D_{\rho|\cdot|^{x}}(\sigma_{1})\times\sigma_{2}+\sigma_{1}\times D_{\rho|\cdot|^{x}}(\sigma_{2}),$
    $D_{\rho|\cdot|^{x}}^{op}(\sigma_{1}\times\sigma_{2})$ $=D_{\rho|\cdot|^{x}}^{op}(\sigma_{1})\times\sigma_{2}+\sigma_{1}\times D_{\rho|\cdot|^{x}}^{op}(\sigma_{2}).$
  2. 2.
    For $a\geq b$,
    $D_{\rho|\cdot|^{x}}(\Delta_{\rho}[a,b])$ =$=\begin{cases}0&\text{ if }x\neq a,\\ \Delta_{\rho}[a-1,b]&\text{ if }x=a.\end{cases}$
    $D_{\rho|\cdot|^{x}}^{op}(\Delta_{\rho}[a,b])$ =$=\begin{cases}0&\text{ if }x\neq b,\\ \Delta_{\rho}[a,b+1]&\text{ if }x=b.\end{cases}$
    $D_{\rho|\cdot|^{x}}(Z_{\rho}[b,a])$ =$=\begin{cases}0&\text{ if }x\neq b,\\ Z_{\rho}[b+1,a]&\text{ if }x=b.\end{cases}$
    $D_{\rho|\cdot|^{x}}^{op}(Z_{\rho}[b,a])$ =$=\begin{cases}0&\text{ if }x\neq a,\\ Z_{\rho}[b,a-1]&\text{ if }x=a.\end{cases}$
  3. 3.
    $D_{\rho|\cdot|^{x}}$ commutes with $D_{\rho|\cdot|^{y}}$ if $|x-y|>1$.
For a multi-set of real numbers $\{x_{1},\dots,x_{r}\}$, we denote the composition of derivatives by
$D_{\rho|\cdot|^{x_{1},\dots,x_{r}}}(\pi):=D_{\rho|\cdot|^{x_{r}}}\circ\cdots\circ D_{\rho|\cdot|^{x_{1}}}(\pi).$
For example, if $\{x,\dots,x\}$ contains $k$ copies of $x$, then
$D_{\rho|\cdot|^{x,\dots,x}}(\pi)=(k!)\cdot D_{\rho|\cdot|^{x}}^{(k)}(\pi).$
Let $\pi$ be a representation of finite length. We define the socle of $\pi$, denoted by $soc(\pi)$, to be the maximal semisimple subrepresentation of $\pi.$
Definition 2.4.
Let $\pi$ be a representation of finite length, $x\in\mathbb{R},$ and $\rho$ be an irreducible unitary self-dual supercuspidal representation of $\mathrm{GL}_{d}(F).$ Define
$S_{\rho|\cdot|^{x}}^{(r)}(\pi):=soc((\rho|\cdot|^{x})^{r}\rtimes\pi).$
For a multi-set of real number $\{x_{1},\dots,x_{r}\}$, we denote the composition of socles by
$S_{\rho|\cdot|^{x_{1},\dots,x_{r}}}(\pi):=S_{\rho|\cdot|^{x_{r}}}\circ\cdots\circ S_{\rho|\cdot|^{x_{1}}}(\pi).$
Theorem 2.5 ([Jan14, Lemma 3.1.3], [AM23, Propositions 3.3, 6.1, Theorem 7.1]).
Let $\rho$ be an irreducible unitary self-dual supercuspidal representation of $\mathrm{GL}_{d}(F),$ $\pi\in\Pi(G_{n}),$ and $x\in\mathbb{R}\setminus\{0\}.$ For any non-negative integers $k$ and $r$, we have the following.
  1. (1)
    The highest $\rho|\cdot|^{x}$-derivative of $\pi$, say $D_{\rho|\cdot|^{x}}^{(k)}(\pi),$ is irreducible.
  2. (2)
    $S_{\rho|\cdot|^{x}}^{(r)}(\pi)$ is irreducible for any $r\geq 0.$
  3. (3)
    $S_{\rho|\cdot|^{x}}^{(k)}(D_{\rho|\cdot|^{x}}^{(k)}(\pi))=\pi,$
    and,
    $D_{\rho|\cdot|^{x}}^{(k+r)}(S_{\rho|\cdot|^{x}}^{(r)}(\pi))=D_{\rho|\cdot|^{x}}^{(k)}(\pi).$
  4. (4)
    The $L$-data of $D_{\rho|\cdot|^{x}}^{(k)}(\pi)$ and $S_{\rho|\cdot|^{x}}^{(k)}(\pi)$ can be explicitly described in terms of those of $\pi.$
When $x=0,$ computing the $\rho$-derivative explicitly is generally difficult. As a remedy, we make use of the $\Delta_{\rho}[0,-1]$-derivative and $Z_{\rho}[0,1]$-derivative, denoted by $D_{\Delta_{\rho}[0,-1]}^{(k)}(\pi)$ and $D_{Z_{\rho}[0,1]}^{(k)}(\pi)$, respectively. These are semisimple representations of $G_{n-2dk}$ defined by
$[Jac_{P_{2dk}}(\pi)]=\Delta_{\rho}[0,-1]^{k}\otimes D_{\Delta_{\rho}[0,-1]}^{(k)}(\pi)+Z_{\rho}[0,1]^{k}\otimes D_{Z_{\rho}[0,1]}^{(k)}(\pi)+\sum_{i}\tau_{i}\otimes\pi_{i},$
where the sum is over all irreducible representations $\tau_{i}$ of $\mathrm{GL}_{2dk}(F)$ such that $\tau_{i}$ is neither isomorphic to $\Delta_{\rho}[0,-1]^{k}$ nor $Z_{\rho}[0,1]^{k}.$ We also define
$S_{\Delta_{\rho}[0,-1]}^{(r)}(\pi):=soc(\Delta_{\rho}[0,-1]^{r}\rtimes\pi),\,\,S_{Z_{\rho}[0,1]}^{(r)}(\pi):=soc(Z_{\rho}[0,1]^{r}\rtimes\pi).$
These derivatives and socles satisfy analogous results as in Theorem 2.5.
Theorem 2.6 ([AM23, Proposition 3.7]).
Let $\rho$ be an irreducible unitary self-dual supercuspidal representation of $\mathrm{GL}_{d}(F)$ and $\pi\in\Pi(G_{n}).$ Assume that $\pi$ is $\rho|\cdot|^{-1}$-reduced (respectively $\rho|\cdot|$-reduced). Then the results of Theorem 2.5(1), (2), and (3) hold with $\rho|\cdot|^{x}$ replaced by $\Delta_{\rho}[0,-1]$ (respectively $Z_{\rho}[0,1]$).
2.3. Local Arthur packets
We recall that a local Arthur parameter is a direct sum of irreducible representations
$\psi:W_{F}\times\mathrm{SL}_{2}(\mathbb{C})\times\mathrm{SL}_{2}(\mathbb{C})\rightarrow\widehat{G}_{n}(\mathbb{C}),$
(2.1) $\psi=\bigoplus_{i=1}^{r}\phi_{i}|\cdot|^{x_{i}}\otimes S_{a_{i}}\otimes S_{b_{i}},$
satisfying the following conditions:
  1. (1)
    $\phi_{i}(W_{F})$ is bounded and consists of semi-simple elements, and $\dim(\phi_{i})=d_{i}$;
  2. (2)
    $x_{i}\in\mathbb{R}$ and $|x_{i}|<\frac{1}{2}$;
  3. (3)
    the restrictions of $\psi$ to the two copies of $\mathrm{SL}_{2}(\mathbb{C})$ are analytic, $S_{k}$ is the $k$-dimensional irreducible representation of $\mathrm{SL}_{2}(\mathbb{C})$, and,
    $\sum_{i=1}^{r}d_{i}a_{i}b_{i}=N:=\begin{cases}2n+1&\text{ when }G_{n}=\mathrm{Sp}_{2n}(F),\\ 2n&\text{ when }G_{n}=\mathrm{SO}_{2n+1}(F).\end{cases}$
We remark that the bound $|x_{i}|<\frac{1}{2}$ follows from the trivial bound of the Ramanujan conjecture of general linear groups.
Two local Arthur parameters are called equivalent if they are conjugate under $\widehat{G}_{n}(\mathbb{C})$. However, in this paper, we do not distinguish $\psi$ with its equivalence class. We let $\Psi^{+}(G_{n})$ be the equivalence class of local Arthur parameters, and $\Psi(G_{n})$ the subset of $\Psi^{+}(G_{n})$ consisting of local Arthur parameters $\psi$ whose restriction to $W_{F}$ is bounded. In other words, $\psi\in\Psi(G_{n})$ if and only if $x_{i}=0$ for $i=1,\dots,r$ in the decomposition (2.1).
By the Local Langlands Correspondence for $\mathrm{GL}_{d_{i}}(F)$, the bounded representations $\phi_{i}$ of $W_{F}$ can be identified with an irreducible unitary supercuspidal representations $\rho_{i}$ of $\mathrm{GL}_{d_{i}}(F)$ ([Hen00, HT01, Sch13]). Consequently, we often write
(2.2) $\psi=\bigoplus_{\rho}\left(\bigoplus_{i\in I_{\rho}}\rho|\cdot|^{x_{i}}\otimes S_{a_{i}}\otimes S_{b_{i}}\right),$
where first sum runs over irreducible unitary supercuspidal representations $\rho$ of $\mathrm{GL}_{n}(F)$, $n\in\mathbb{Z}_{\geq 1}$. Occasionally, we also write $\rho|\cdot|^{x}\otimes S_{a}=\rho|\cdot|^{x}\otimes S_{a}\otimes S_{1}.$
Let $\psi$ be a local Arthur parameter as in (2.2), we say that $\psi$ is of good parity if $\psi\in\Psi(G_{n})$ (i.e., $x_{i}=0$ for all $i$) and every summand $\rho\otimes S_{a_{i}}\otimes S_{b_{i}}$ is self-dual and of the same type as $\psi.$ That is, $\rho$ is self dual and
  • if $G_{n}=\mathrm{Sp}_{2n}(F)$ and $\rho$ is orthogonal (resp. symplectic), then $a_{i}+b_{i}$ is even (resp. odd);
  • if $G_{n}=\mathrm{SO}_{2n+1}(F)$ and $\rho$ is orthogonal (resp. symplectic), then $a_{i}+b_{i}$ is odd, (resp. even).
We let $\Psi_{gp}(G_{n})$ denote the subset of $\Psi(G_{n})$ consisting of local Arthur parameters of good parity.
Let $\psi\in\Psi^{+}(G_{n}).$ From the decomposition (2.2), we define a subrepresentation $\psi_{nu,>0}$ of $\psi$ by
$\psi_{nu,>0}:=\bigoplus_{\rho}\left(\bigoplus_{\begin{subarray}{c}i\in I_{\rho},\\ x_{i}>0\end{subarray}}\rho|\cdot|^{x_{i}}\otimes S_{a_{i}}\otimes S_{b_{i}}\right).$
Since the image of $\psi$ is contained in $\widehat{G}_{n}(\mathbb{C})$, $\psi$ is self-dual, and hence $\psi$ also contains $(\psi_{nu,>0})^{\vee}$. We define $\psi_{u}\in\Psi(G_{m})$ for some $m\leq n$ by
(2.3) $\psi=\psi_{nu,>0}\oplus\psi_{u}\oplus(\psi_{nu,>0})^{\vee}.$
Equivalently, we have
$\psi_{u}:=\bigoplus_{\rho}\left(\bigoplus_{\begin{subarray}{c}i\in I_{\rho},\\ x_{i}=0\end{subarray}}\rho\otimes S_{a_{i}}\otimes S_{b_{i}}\right).$
In [Art13, Theorem 1.5.1], for a local Arthur parameter $\psi\in\Psi(G_{n})$, Arthur constructed a finite multi-set $\Pi_{\psi}$ consisting of irreducible unitary representations of $G_{n}.$ We call $\Pi_{\psi}$ the local Arthur packet of $\psi.$ Mœglin gave another construction of $\Pi_{\psi}$ and showed that it is multiplicity-free ([Mœ11a]). For $\psi\in\Psi^{+}(G_{n})$, Arthur defined the local Arthur packet $\Pi_{\psi}$ ([Art13, (1.5.1)]), by
(2.4) $\Pi_{\psi}:=\{\tau_{\psi_{nu,>0}}\rtimes\pi_{u}\ |\ \pi_{u}\in\Pi_{\psi_{u}}\},$
where $\tau_{\psi_{nu,>0}}$ is the following irreducible representation of a general linear group
$\tau_{\psi_{nu,>0}}=\bigtimes_{\rho}\bigtimes_{i\in I_{\rho}}\begin{pmatrix}\frac{a_{i}-b_{i}}{2}+x_{i}&\cdots&\frac{a_{i}+b_{i}}{2}-1+x_{i}\\ \vdots&\ddots&\vdots\\ \frac{-a_{i}-b_{i}}{2}+1+x_{i}&\cdots&\frac{b_{i}-a_{i}}{2}+x_{i}\end{pmatrix}_{\rho}.$
Since $|x_{i}|<\frac{1}{2}$ in the decomposition (2.2), the parabolic induction in (2.4) is always irreducible by [Mœ11b, Proposition 5.1] (see also [Jan97, Theorem 9.3(6)], [Tad09, Proposition 3.2(i)]). We say that an irreducible representation $\pi$ of $G_{n}$ is of Arthur type if $\pi\in\Pi_{\psi}$ for some local Arthur parameter $\psi\in\Psi^{+}(G_{n})$.
Next, we further decompose $\psi_{u}$. Suppose $\rho\otimes S_{a}\otimes S_{b}$ is an irreducible summand of $\psi_{u}$ that is either not self-dual, or self-dual but not of the same type as $\psi$. Then $\psi$ must also contain the another summand $(\rho\otimes S_{a}\otimes S_{b})^{\vee}=\rho^{\vee}\otimes S_{a}\otimes S_{b}$. Therefore, we may choose a subrepresentation $\psi_{np}$ of $\psi_{u}$ such that
(2.5) $\psi_{u}=\psi_{np}\oplus\psi_{gp}\oplus\psi_{np}^{\vee},$
where $\psi_{gp}$ is of good parity, and any irreducible summand of $\psi_{np}$ is either not self-dual or self-dual but not of the same type as $\psi$. In [Mœ06a], Mœglin constructed the local Arthur packet $\Pi_{\psi_{u}}$ from $\Pi_{\psi_{gp}}$, which we record below.
Theorem 2.7 ([Mœ06a, Theorem 6], [Xu17b, Proposition 8.11]).
Let $\psi_{u}\in\Psi(G_{n})$ with a choice of decomposition (2.5). Write
$\psi_{np}=\bigoplus_{\rho}\left(\bigoplus_{i\in I_{\rho}}\rho\otimes S_{a_{i}}\otimes S_{b_{i}}\right),$
and consider the following irreducible parabolic induction
$\tau_{\psi_{np}}=\bigtimes_{\rho}\bigtimes_{i\in I_{\rho}}\begin{pmatrix}\frac{a_{i}-b_{i}}{2}&\cdots&\frac{a_{i}+b_{i}}{2}-1\\ \vdots&\ddots&\vdots\\ \frac{-a_{i}-b_{i}}{2}+1&\cdots&\frac{b_{i}-a_{i}}{2}\end{pmatrix}_{\rho}.$
Then for any $\pi_{gp}\in\Pi_{\psi_{gp}}$ the induced representation $\tau_{\psi_{np}}\rtimes\pi_{gp}$ is irreducible, independent of choice of $\psi_{np}$. Moreover,
$\Pi_{\psi}=\{\tau_{\psi_{np}}\rtimes\pi_{gp}\,|\,\pi\in\Pi_{\psi_{gp}}\}.$
Combined with (2.4), we obtain the following.
Theorem 2.8 ([Mœ11b, Proposition 5.1]).
Let $\psi\in\Psi^{+}(G_{n})$ with decomposition $\psi=\psi_{nu,>0}+\psi_{np}+\psi_{gp}+\psi_{np}^{\vee}+\psi_{nu,>0}^{\vee}$ as above. Then, for any $\pi_{gp}\in\Pi_{\psi_{gp}},$ the induction $\tau_{\psi_{nu,>0}}\times\tau_{\psi_{np}}\rtimes\pi_{gp}$ is irreducible. As a consequence,
(2.6) $\Pi_{\psi}=\{\tau_{\psi_{nu,>0}}\times\tau_{\psi_{np}}\rtimes\pi_{gp},|,\pi_{gp}\in\Pi_{\psi_{gp}}\}.$
Theorem 2.8 reduces calculating intersections of local Arthur packets to the good parity case. We give an analogous definition of good parity for representations.
Definition 2.9.
We say an irreducible representation
$\pi=L(\Delta_{\rho_{1}}[x_{1},y_{1}],\dots,\Delta_{\rho_{r}}[x_{r},y_{r}];\pi_{temp})$
of $G_{n}$ is of good parity if the following hold:
  1. $\bullet$
    The tempered representation $\pi_{temp}$ lies in $\Pi_{\psi_{temp}}$ for some $\psi_{temp}\in\Psi_{gp}(G_{m})$.
  2. $\bullet$
    For $1\leq i\leq r$, $x_{i},y_{i}\in\frac{1}{2}\mathbb{Z}$ and $\rho_{i}\otimes S_{x_{i}-y_{i}+1}\otimes S_{1}$ is self-dual of the same type as $\widehat{G}_{n}$.
Let $\pi\in\Pi_{\psi}$. By the construction of local Arthur packets in the good parity case and Theorem 2.7, we have that $\pi$ is of good parity if and only if $\psi$ is of good parity. With Theorem 2.8, we reduce the computation of $\Psi(\pi)$ to $\Psi(\pi_{gp})$.
Corollary 2.10 ([HLL22, Corollary 2.13(1)]).
For $\pi\in\Pi(G_{n})$, there exists $\psi\in\Psi^{+}(G_{n})$ such that $\pi\in\Pi_{\psi}$ if and only if $\pi$ is of the form
$\pi=\tau_{\psi_{nu,>0}}\times\tau_{\psi_{np}}\rtimes\pi_{gp},$
and $\pi_{gp}\in\Pi_{\psi_{gp}}$ for some $\psi_{gp}\in\Psi_{gp}(G_{n})$. Moreover, we have
$\Psi(\pi):=$$\{\psi\in\Psi^{+}(G_{n})\ |\ \pi\in\Pi_{\psi}\}$
$=$ $\{\psi_{nu,>0}+\psi_{np}+\psi_{gp}+\psi_{np}^{\vee}+\psi_{nu,>0}\ |\ \psi_{gp}\in\Psi(\pi_{gp})\}.$
2.4. Parametrization of tempered spectrum
Let $\psi\in\Psi(G_{n})$. An important ingredient of the construction of local Arthur packets is a map
(2.7) $\Pi_{\psi}$ $\longrightarrow\widehat{\mathcal{S}}_{\psi},$
$\pi$ $\longmapsto\langle\cdot,\pi\rangle,$
where $\widehat{\mathcal{S}}_{\psi}$ denotes Pontryagin dual of the component group
$\mathcal{S}_{\psi}:=\pi_{0}(\mathrm{Cent}_{\widehat{G}_{n}(\mathbb{C})}(\mathrm{Im}(\psi))/Z(\widehat{G}_{n}(\mathbb{C}))).$
The map (2.7) is neither injective nor surjective in general. However, when $\psi$ is tempered, it is a bijection (see Theorem 2.11 below). In this subsection, we recall the combinatorial description of $\widehat{\mathcal{S}}_{\psi}$ in [Xu17b, §2] and use it to give a parametrization of the tempered spectrum of $G_{n}$.
Write $\psi=\psi_{np}+\psi_{gp}+\psi_{np}^{\vee}$. There is a bijection between $\mathcal{S}_{\psi}$ and $\mathcal{S}_{\psi_{gp}}$. Therefore, to describe $\widehat{\mathcal{S}}_{\psi}$, it suffices to assume $\psi\in\Psi_{gp}(G_{n})$. Write
(2.8) $\psi=\bigoplus_{\rho}\bigoplus_{i\in I_{\rho}}\rho\otimes S_{a_{i}}\otimes S_{b_{i}}.$
First, we consider the enhanced component group of $\psi$ defined by
$\mathcal{A}_{\psi}=\bigoplus_{\rho}\bigoplus_{i\in I_{\rho}}(\mathbb{Z}/2\mathbb{Z})\alpha_{\rho,i}.$
Here, $\mathcal{A}_{\psi}$ is the finite vector space over $\mathbb{Z}/2\mathbb{Z}$ with basis $\alpha_{\rho,i}$ corresponding to the summands $\rho\otimes S_{a_{i}}\otimes S_{b_{i}}$ of Equation (2.8). It is possible that for some $i,j\in I_{\rho}$, we have $\rho\otimes S_{a_{i}}\otimes S_{b_{i}}=\rho\otimes S_{a_{j}}\otimes S_{b_{j}}$; however, we distinguish these summands in $\mathcal{A}_{\psi},$ i.e., $\alpha_{\rho,i}\neq\alpha_{\rho,j}$ in $\mathcal{A}_{\psi}.$ The central element of $\mathcal{A}_{\psi}$ is $z_{\psi}:=\sum_{\rho}\sum_{i\in I_{\rho}}\alpha_{\rho,i}.$
The component group $\mathcal{S}_{\psi}$ of $\psi$ can be identified with the quotient of $\mathcal{A}_{\psi}$ by the subgroup generated by the central element and the elements $\alpha_{\rho,i}+\alpha_{\rho,j}$ such that $i,j\in I_{\rho}$ with $\rho\otimes S_{a_{i}}\otimes S_{b_{i}}=\rho\otimes S_{a_{j}}\otimes S_{b_{j}}.$ Hence, we identify $\widehat{\mathcal{S}}_{\psi}$ as the set of functions $\varepsilon$ from the irreducible summands of $\psi$ to {±$\{\pm 1\}$ that satisfy
  1. $\bullet$
    $\varepsilon(\rho\otimes S_{a_{i}}\otimes S_{b_{i}})=\varepsilon(\rho\otimes S_{a_{j}}\otimes S_{b_{j}})$ if $\rho\otimes S_{a_{i}}\otimes S_{b_{i}}=\rho\otimes S_{a_{i}}\otimes S_{b_{i}}$, and
  2. $\bullet$
    $\prod_{\rho}\prod_{i\in I_{\rho}}\varepsilon(\rho\otimes S_{a_{i}}\otimes S_{b_{i}})=1$.
A local Arthur parameter $\psi,$ decomposed as in Equation (2.8), is tempered if $b_{i}=1$ for any $\rho$ and $i\in I_{\rho},$ i.e., $\psi$ is trivial on the second $\mathrm{SL}_{2}(\mathbb{C}).$ We say that a local Arthur packet $\Pi_{\psi}$ is tempered if $\psi$ is tempered. A Whittaker datum for $G_{n}$ is a $G_{n}$-conjugacy class of a tuple $(B,\chi)$, where $B$ is an $F$-rational Borel subgroup of $\mathrm{G}_{n}$ and $\chi$ is a generic character of the $F$-points of the unipotent radical of $B$. The following theorem is Arthur’s classification of the tempered representations.
Theorem 2.11 ([Art13, Theorem 1.5.1]).
Any irreducible tempered representation of $G_{n}$ lies in $\Pi_{\psi}$ for some tempered local Arthur parameter $\psi.$ Moreover, if $\psi_{1}$ and $\psi_{2}$ are two non-isomorphic tempered local Arthur parameters, then
$\Pi_{\psi_{1}}\cap\Pi_{\psi_{2}}=\emptyset.$
Finally, if one fixes a choice of Whittaker datum for $G_{n}$ and $\psi$ is tempered, then there is a bijective map between the tempered local Arthur packet $\Pi_{\psi}$ and $\widehat{\mathcal{S}}_{\psi}.$
Hereinafter, we fix a choice of Whittaker datum for $G_{n}.$ When $\psi$ is tempered and of good parity, we write $\pi(\psi,\varepsilon)$ for the representation in $\Pi_{\psi}$ corresponding to $\varepsilon\in\widehat{\mathcal{S}}_{\psi}$ via the bijection in Theorem 2.11.
Following the notation in [AM23], when
$\psi=\bigoplus_{i=1}^{m}\rho_{i}\otimes S_{a_{i}}\otimes S_{1}$
is a tempered local Arthur parameter, we set $x_{i}=\frac{a_{i}-1}{2}$ for $i=1,\dots,m$. We also express $\pi(\psi,\varepsilon)\in\Pi_{\psi}$, the representation corresponding to $\varepsilon$ via Theorem 2.11, as
$\pi((x_{1},\rho_{1})^{\varepsilon(\rho_{1}\otimes S_{a_{1}})},\dots,(x_{m},\rho_{m})^{\varepsilon(\rho\otimes S_{a_{m}})}).$
When $\psi$ has only one $\rho$ in its decomposition, we often ignore it in the notation and simply write
$\pi(x_{1}^{\varepsilon(\rho\otimes S_{a_{1}})},\dots,x_{m}^{\varepsilon(\rho\otimes S_{a_{m}})})=\pi(\psi,\varepsilon).$
3. Atobe’s reformulation
In this section, we recall the main definitions and results of [Ato22b] for the construction of local Arthur packets of good parity. We also recall the operators of [HLL22] and various results, including an algorithm to determine when a representation is of Arthur type (see Algorithm 3.9).
We fix the following notation throughout this section. Let $\psi$ be any local Arthur parameter of good parity with decomposition
$\psi=\bigoplus_{\rho}\bigoplus_{i\in I_{\rho}}\rho\otimes S_{a_{i}}\otimes S_{b_{i}}.$
We set $A_{i}=\frac{a_{i}+b_{i}}{2}-1$ and $B_{i}=\frac{a_{i}-b_{i}}{2}$ for $i\in I_{\rho}.$
A total order $>_{\psi}$ on $I_{\rho}$ is called admissible if satisfies:
For $i,j\in I_{\rho}$, if $A_{i}>A_{j}$ and $B_{i}>B_{j}$, then $i>_{\psi}j$.
Sometimes we consider an order $>_{\psi}$ on $I_{\rho}$ satisfying:
For $i,j\in I_{\rho}$, if $B_{i}>B_{j}$, then $i>_{\psi}j$.
Note that ($P^{\prime}$) implies ($P$). Often, we write $>$ instead of $>_{\psi}$ when the admissible order is fixed.
Suppose now that we have fixed an admissible order for $\psi.$ We define the collection of ordered multi-sets
$\operatorname{supp}(\psi):=\cup_{\rho}\{[A_{i},B_{i}]_{\rho}\}_{i\in(I_{\rho},>)}.$
This is called the support of $\psi$. Note that $\operatorname{supp}(\psi)$ depends implicitly on the fixed admissible order.
3.1. Extended multi-segments and associated representations
In this subsection, we recall Atobe’s parameterization of local Arthur packets from [Ato22b].
Definition 3.1 ([Ato22a, Definition 3.1]).
(Extended multi-segments)
  1. (1)
    An extended segment is a triple $([A,B]_{\rho},l,\eta)$, where
    • $[A,B]_{\rho}=\{\rho|\cdot|^{A},\rho|\cdot|^{A-1},\dots,\rho|\cdot|^{B}\}$ is a segment for an irreducible unitary supercuspidal representation $\rho$ of some $\mathrm{GL}_{d}(F)$;
    • $l\in\mathbb{Z}$ with $0\leq l\leq\frac{b}{2}$, where $b=\#[A,B]_{\rho}=A-B+1$;
    • $\eta\in\{\pm 1\}$.
  2. (2)
    An extended multi-segment for $G_{n}$ is an equivalence class (via the equivalence defined below) of multi-sets of extended segments
    $\mathcal{E}=\cup_{\rho}\{([A_{i},B_{i}]_{\rho},l_{i},\eta_{i})\}_{i\in(I_{\rho},>)}$
    such that
    • $I_{\rho}$ is a totally ordered finite set with a fixed admissible total order $>$;
    • $A_{i}+B_{i}\geq 0$ for all $\rho$ and $i\in I_{\rho}$;
    • as a representation of $W_{F}\times\mathrm{SL}_{2}(\mathbb{C})\times\mathrm{SL}_{2}(\mathbb{C})$,
      $\psi_{\mathcal{E}}=\bigoplus_{\rho}\bigoplus_{i\in I_{\rho}}\rho\otimes S_{a_{i}}\otimes S_{b_{i}}$
      where $(a_{i},b_{i})=(A_{i}+B_{i}+1,A_{i}-B_{i}+1)$, is a local Arthur parameter for $G_{n}$ of good parity. We shall denote $\psi_{\mathcal{E}}$ the local Arthur parameter associated with $\mathcal{E}$.
    • The sign condition
      (3.1) $\prod_{\rho}\prod_{i\in I_{\rho}}(-1)^{[\frac{b_{i}}{2}]+l_{i}}\eta_{i}^{b_{i}}=1$
      holds.
  3. (3)
    Two extended segments $([A,B]_{\rho},l,\eta)$ and $([A^{\prime},B^{\prime}]_{\rho^{\prime}},l^{\prime},\eta^{\prime})$ are weakly equivalent if
    • $[A,B]_{\rho}=[A^{\prime},B^{\prime}]_{\rho^{\prime}}$;
    • $l=l^{\prime}$; and
    • $\eta=\eta^{\prime}$ whenever $l=l^{\prime}<\frac{b}{2}$.
    Two extended multi-segments $\mathcal{E}=\cup_{\rho}\{([A_{i},B_{i}]_{\rho},l_{i},\eta_{i})\}_{i\in(I_{\rho},>)}$ and $\mathcal{E}^{\prime}=\cup_{\rho}\{([A^{\prime}_{i},B^{\prime}_{i}]_{\rho},l^{\prime}_{i},\eta^{\prime}_{i})\}_{i\in(I_{\rho},>)}$ are weakly equivalent if for any $\rho$ and $i\in I_{\rho}$, the extended segments $([A_{i},B_{i}]_{\rho},l_{i},\eta_{i})$ and $([A^{\prime}_{i},B^{\prime}_{i}]_{\rho},l^{\prime}_{i},\eta^{\prime}_{i})$ are weakly equivalent.
  4. (4)
    We define the support of $\mathcal{E}$ to be the collection of ordered multi-sets
    $\operatorname{supp}(\mathcal{E})=\cup_{\rho}\{[A_{i},B_{i}]_{\rho}\}_{i\in(I_{\rho},>)}.$
If the admissible order $>$ is clear in the context, for $k\in I_{\rho}$, we often denote $k+1\in I_{\rho}$ to be the unique element adjacent with $k$ and $k+1>k$.
We attach a pictograph to each extended multi-segment by the same way in [Ato22b, §3]. Here is an example.
Example 3.2 ([HLL22, Example 3.2]).
Let $\rho$ be the trivial representation. The pictograph
[…]
corresponds to the extended multi-segment =$\mathcal{E}=\{([A_{i},B_{i}]_{\rho},l_{i},\eta_{i})\}_{i\in(1<2<3)}$ of $\mathrm{Sp}_{44}(F)$ where
  1. $\bullet$
    $([A_{1},B_{1}]_{\rho},[A_{2},B_{2}]_{\rho},[A_{3},B_{3}]_{\rho})=([4,-1]_{\rho},[3,2]_{\rho},[4,4]_{\rho})$ specify the “support” of each row.
  2. $\bullet$
    $(l_{1},l_{2},l_{3})=(2,1,0)$ counts the number of pairs of triangles in each row.
  3. $\bullet$
    $(\eta_{1},\eta_{2},\eta_{3})=(1,1,-1)$ records the sign of the first circle in each row. Note that $\eta_{2}=1,-1$ are weakly equivalent.
The associated local Arthur parameter is
$\psi_{\mathcal{E}}=\rho\otimes S_{4}\otimes S_{6}+\rho\otimes S_{6}\otimes S_{2}+\rho\otimes S_{9}\otimes S_{1}.$
For each extended multi-segment $\mathcal{E}$, Atobe associated a representation $\pi(\mathcal{E})$, which is irreducible or zero, see [Ato22b, §3.2]. The $L$-data of $\pi(\mathcal{E})$ can be computed explicitly by the algorithms in [AM23]. This leads to Atobe’s parametrization of the local Arthur packets associated with local Arthur parameters of good parity.
Theorem 3.3 ([Ato22b, Theorem 3.3]).
Suppose $\psi=\bigoplus_{\rho}\bigoplus_{i\in I_{\rho}}\rho\otimes S_{a_{i}}\otimes S_{b_{i}}$ is a local Arthur parameter of good parity of $G_{n}$. Choose an admissible order $>_{\psi}$ on $I_{\rho}$ for each $\rho$ that satisfies ($P^{\prime}$) if $\frac{a_{i}-b_{i}}{2}<0$ for some $i\in I_{\rho}$. Then
$\bigoplus_{\pi\in\Pi_{\psi}}\pi=\bigoplus_{\mathcal{E}}\pi(\mathcal{E}),$
where $\mathcal{E}$ runs over all extended multi-segments with $\operatorname{supp}(\mathcal{E})=\operatorname{supp}(\psi)$ and $\pi(\mathcal{E})\neq 0$.
By $\operatorname{supp}(\mathcal{E})=\operatorname{supp}(\psi),$ we mean that they are equal as multi-sets and also that the admissible orders on $I_{\rho}$ agree.
Often, we need to study an extended multi-segment $\mathcal{E}$ piece by piece. Suppose
$\mathcal{E}=\cup_{\rho}\{([A_{i},B_{i}]_{\rho},l_{i},\eta_{i})\}_{i\in(I_{\rho},>)}$
satisfies $\pi(\mathcal{E})\neq 0$. We set
$\mathcal{E}_{\rho}=\{([A_{i},B_{i}]_{\rho},l_{i},\eta_{i})\}_{i\in(I_{\rho},>)},\ \mathcal{E}^{\rho}=\cup_{\rho^{\prime}\not\cong\rho}\{([A_{i},B_{i}]_{\rho^{\prime}},l_{i},\eta_{i})\}_{i\in(I_{\rho^{\prime}},>)}.$
For $i\in I_{\rho}$, we usually call the extended segment $([A_{i},B_{i}]_{\rho},l_{i},\eta_{i})$ the $i$-th row of $\mathcal{E}_{\rho}$. Here are the sets of extended multi-segments that we are interested in the most.
Definition 3.4.
We define $\underline{\mathrm{Rep}}$ to be the set of extended multi-segments $\mathcal{E}$ such that $\pi(\mathcal{E})\neq 0$. Furthermore, we let $\underline{\mathrm{Rep}}^{(P^{\prime})}$ be the subset of $\underline{\mathrm{Rep}}$ consisting of extended multi-segments whose admissible orders already satisfy ($P^{\prime}$).
3.2. Intersection of local Arthur packets
In this subsection, we recall the main results in [HLL22] and [Ato23] on the intersection of local Arthur packets. We shall state them with the notation in [HLL22].
The main ingredients of these results are operators $R_{k}$, $ui_{i,j}$, $dual,$ $dual_{i}^{\pm}$ on extended multi-segments. In this paper, we only need certain consequences of the definition of $ui_{i,j}$ and $dual_{i}^{\pm}$. We state these consequences in Lemma 4.1 and refer the explicit definitions of $ui_{i,j},dual_{i}^{\pm}$ to [HLL22, Definition 3.23, 5.1, 6.5]. On the other hand, we recall the definition $dual$ and its compatibility with Aubert-Zelevinsky dual now.
Let $\widehat{\pi}$ denote the Aubert-Zelevinsky dual of $\pi$ defined in [Aub95]. The Aubert-Zelevinsky dual is compatible with local Arthur packets in the following sense. Let $\psi=\bigoplus_{\rho}\bigoplus_{i\in I_{\rho}}\rho\otimes S_{a_{i}}\otimes S_{b_{i}}$ be a local Arthur parameter of good parity and set $\widehat{\psi}:=\bigoplus_{\rho}\bigoplus_{i\in I_{\rho}}\rho\otimes S_{b_{i}}\otimes S_{a_{i}}.$ Note that $\widehat{\psi}$ is also a local Arthur parameter of good parity. By the results in [Xu17b, §A], we have
$\Pi_{\widehat{\psi}}=\{\widehat{\pi}\,|\,\pi\in\Pi_{\psi}\}.$
Therefore, given $\mathcal{E}\in\underline{\mathrm{Rep}}$, it is natural to ask what is the formula for the extended multi-segment $dual(\mathcal{E})$ such that $\pi(dual(\mathcal{E}))\cong\widehat{\pi(\mathcal{E})}$ and $\psi_{\mathcal{E}^{\prime}}=\widehat{\psi}_{\mathcal{E}}$. This is answered in [Ato22b, §6].
Definition 3.5 ([Ato22b, Definition 6.1]).
Let $\mathcal{E}=\cup_{\rho}\{([A_{i},B_{i}]_{\rho},l_{i},\eta_{i})\}_{i\in(I_{\rho},>)}$ be an extended multi-segment such that the admissible order $>$ on $I_{\rho}$ satisfies (P’) for all $\rho$. We define
$dual(\mathcal{E})=\cup_{\rho}\{([A_{i},-B_{i}]_{\rho},l_{i}^{\prime},\eta_{i}^{\prime})\}_{i\in(I_{\rho},>^{\prime})}$
as follows:
  1. (1)
    The order $>^{\prime}$ is defined by $i>^{\prime}j$ if and only if $j>i.$
  2. (2)
    We set
    $l_{i}^{\prime}=\begin{cases}l_{i}+B_{i}&\mathrm{if}\,B_{i}\in\mathbb{Z},\\ l_{i}+B_{i}+\frac{1}{2}(-1)^{\alpha_{i}}\eta_{i}&\mathrm{if}\,B_{i}\not\in\mathbb{Z},\end{cases}$
    and
    $\eta_{i}^{\prime}=\begin{cases}(-1)^{\alpha_{i}+\beta_{i}}\eta_{i}&\mathrm{if}\,B_{i}\in\mathbb{Z},\\ (-1)^{\alpha_{i}+\beta_{i}+1}\eta_{i}&\mathrm{if}\,B_{i}\not\in\mathbb{Z},\end{cases}$
    where $\alpha_{i}=\sum_{j\in I_{\rho},j<i}a_{j},$ and $\beta_{i}=\sum_{j\in I_{\rho},j>i}b_{j},$ $a_{j}=A_{j}+B_{j}+1$, $b_{j}=A_{j}-B_{j}+1$.
  3. (3)
    When $B_{i}\not\in\mathbb{Z}$ and $l_{i}=\frac{b_{i}}{2}$, we set $\eta_{i}=(-1)^{\alpha_{i}+1}.$
If $\mathcal{F}=\mathcal{E}_{\rho}$, we define $dual(\mathcal{F}):=(dual(\mathcal{E}))_{\rho}$.
Now we state the main theorems of [HLL22].
Theorem 3.6 ([HLL22, Theorem 1.4]).
Let $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ be two extended multi-segments for $G_{n}$. Suppose that $\pi(\mathcal{E}_{1})\in\Pi_{\psi}$. Then the followings hold.
  1. (1)
    Let $T$ be any of the four operators $R_{k}$, $ui_{i,j}$, $dual\circ ui_{i,j}\circ dual$, $dual_{k}$ or their inverses. We have $T(\mathcal{E}_{1})$ is also an extended multi-segment for $G_{n}$, and
    $\pi(T(\mathcal{E}_{1}))=\pi(\mathcal{E}_{1}).$
  2. (2)
    We have $\pi(\mathcal{E}_{1})\cong\pi(\mathcal{E}_{2})$ if and only if $\mathcal{E}_{2}$ can be obtained from $\mathcal{E}_{1}$ by a finite chain of the four operators and their inverses in Part (1).
  3. (3)
    There is a precise formula/algorithm to compute the set $\{\psi^{\prime}\ |\ \pi(\mathcal{E}_{1})\in\Pi_{\psi^{\prime}}\}$.
We remark that the operator $R_{k}$ does not change the corresponding local Arthur parameter, i.e. $\psi_{\mathcal{E}}=\psi_{R_{k}(\mathcal{E})}$. Thus, we often omit $R_{k}$ in the computation for $\Psi(\pi)$.
3.3. An algorithm to determine Arthur type
In this subsection, we review an algorithm ([HLL22, Algorithm 7.9], see Algorithm 3.9 below) which determines if a given representation is of Arthur type. We begin by recalling some notation needed for the algorithm, along with some results which are helpful to us later.
Definition 3.7 ([HLL22, Definition 4.1]).
  1. (1)
    Suppose
    =$\mathcal{F}=\{([A_{i},B_{i}]_{\rho},l_{i},\eta_{i})\}_{i\in(I_{\rho,>})}.$
    For example, let $\mathcal{F}=\mathcal{E}_{\rho}$ for some extended multi-segment $\mathcal{E}$. We define ordered multi-sets
    $\Omega(\mathcal{F}):=$ $\sum_{i\in I_{\rho}}[A_{i},B_{i}]_{\rho}=\{\rho|\cdot|^{\alpha_{1}},\dots,\rho|\cdot|^{\alpha_{t}}\},$
    $\overline{\Omega(\mathcal{F})}:=$ $\sum_{i\in I_{\rho}}[B_{i},-A_{i}]_{\rho}=\{\rho|\cdot|^{\beta_{1}},\dots,\rho|\cdot|^{\beta_{r}}\},$
    where $\alpha_{1}\leq\cdots\leq\alpha_{t}$ and $\beta_{1}\geq\cdots\geq\beta_{r}$. Suppose $\mathcal{E}=\cup_{\rho}\mathcal{E}_{\rho}$ is an extended multi-segment. Fix an arbitrary order on the set
    $\{\rho\ |\ \mathcal{E}_{\rho}\neq\emptyset\}=\{\rho_{1},\dots,\rho_{r}\}.$
    We define multi-sets $\Omega(\mathcal{E})$ and $\overline{\Omega(\mathcal{E})}$ to be the sum of multi-sets
    $\Omega(\mathcal{E})$ $:=\Omega(\mathcal{E}_{\rho_{1}})+\cdots+\Omega(\mathcal{E}_{\rho_{r}}),$
    $\overline{\Omega(\mathcal{E})}$ $:=\overline{\Omega(\mathcal{E}_{\rho_{1}})}+\cdots+\overline{\Omega(\mathcal{E}_{\rho_{r}})}.$
  2. (2)
    For each ordered multi-set $\Omega=\{\rho_{1}|\cdot|^{\gamma_{1}},\dots,\rho_{t}|\cdot|^{\gamma_{t}}\}$, we define
    $D_{\Omega}$ $:=D_{\rho_{t}|\cdot|^{\gamma_{t}}}\circ\cdots\circ D_{\rho_{1}|\cdot|^{\gamma_{1}}},$
    $S_{\Omega}$ $:=S_{\rho_{1}|\cdot|^{\gamma_{1}}}\circ\cdots\circ S_{\rho_{t}|\cdot|^{\gamma_{t}}}.$
    For an extended multi-segment $\mathcal{E}$, we define $D_{\Omega(\mathcal{E})}=\circ_{\rho}D_{\Omega(\mathcal{E}_{\rho})}$. Note that the derivative is independent of the composition order of $D_{\Omega(\mathcal{E}_{\rho})}$.
Similarly to Definition 3.7, we define a multi-set attached to a representation.
Definition 3.8 ([HLL22, Definition 4.6]).
For an irreducible representation
$\pi=L\left(\Delta_{\rho_{1}}[x_{1},-y_{1}],\dots,\Delta_{\rho_{t}}[x_{t},-y_{t}];\pi(\sum_{j=t+1}^{m}\rho_{j}\otimes S_{2z_{j}+1},\varepsilon)\right),$
We define
$\Omega(\pi):=\{\rho_{1}|\cdot|^{x_{1}},\dots,\rho_{t}|\cdot|^{x_{t}}\}+\{\rho_{1}|\cdot|^{y_{1}},\dots,\rho_{t}|\cdot|^{y_{t}}\}+\{\rho_{t+1}|\cdot|^{z_{t+1}},\dots,\rho_{m}|\cdot|^{z_{m}}\}.$
We denote $\Omega(\pi)_{\rho}$ to be the maximal sub-multi-set of $\Omega(\pi)$ whose elements are all of the form $\rho|\cdot|^{x}$ for some $x\in\mathbb{R}$.
We now state an algorithm to determine when a representation is of Arthur type.
Algorithm 3.9 ([HLL22, Algorithm 7.9]).
Given a representation $\pi$ of good parity, proceed as follows:
  1. Step 0:
    Set $\psi=0$. Repeat steps 1 to 4 for each $\rho\in\{\rho\ |\ \Omega(\pi)_{\rho}\neq\emptyset\}$.
  2. Step 1:
    Let $A=\max\{x\ |\ \rho|\cdot|^{x}\in\Omega(\pi)_{\rho}\}$ and $\epsilon\in\{0,1/2\}$ such that $A+\epsilon\in\mathbb{Z}$. Set $\Omega^{+}=\emptyset=\Omega^{-}$.
  3. Step 2:
    Compute the following set
    $\mathcal{B}=$$\{B>1/2\ |\ D_{\rho|\cdot|^{B}}(\pi)\neq 0\}=\{B_{1},\dots,B_{r}\},$
    where $B_{i}$ is decreasing. For each $1\leq i\leq r$, we compute (by [AM23, Theorem 7.1]) recursively the integer $k_{i,t}$ and representation $\pi_{i,t}$ for $t$ in the segment $[A+1,B_{i}-1]$ via
    $\begin{cases}\pi_{i,B_{i}-1}=\pi,\\ \pi_{i,t}=D_{\rho|\cdot|^{t}}^{(k_{i,t})}(\pi_{i,t-1})\text{ is the highest derivative.}\end{cases}$
    If $k_{i,A+1}\neq 0$, then $\pi$ is not of Arthur type and the procedure ends. Set $k_{i,t}:=0$ if $t\not\in[A+1,B_{i}-1]$.
    Denote $K_{i,t}:=k_{i,t}-k_{i-1,t}$. For $t\in[A+1,B_{i}+1]$, if $K_{i,t}>K_{i,t-1}$, then $\pi$ is not of Arthur type and the procedure ends. If $K_{i,t}<K_{i,t-1}$, then add $K_{i,t-1}-K_{i,t}$ copies of $\rho\otimes S_{(t-1)+B_{i}+1}\otimes S_{(t-1)-B_{i}+1}$ to $\psi$ and add the same copies of elements in the segment $[t-1,B_{i}]_{\rho}$ in $\Omega^{+}$.
  4. Step 3:
    Reorder
    $\Omega^{+}=\{\rho|\cdot|^{x_{1}},\dots,\rho|\cdot|^{x_{r}}\}$
    such that $x_{1}\leq\cdots\leq x_{r}$. For $t\in[A+\epsilon,1]$, denote
    $sh^{t}(\Omega^{+})=\{\rho|\cdot|^{x_{1}+t},\dots,\rho|\cdot|^{x_{r}+t}\}.$
    Compute the representation (by [AM23, Theorem 7.1])
    $\pi_{A}=S_{sh^{A+\epsilon}(\Omega^{+})}\circ\cdots\circ S_{sh^{1}(\Omega^{+})}(\pi),$
    and the set
    ¯=$\overline{\mathcal{B}}=\{B<0\ |\ D_{\rho|\cdot|^{B}}(\pi_{A})\neq 0\}=\{\overline{B_{1}},\dots,\overline{B_{\overline{r}}}\},$
    where $\overline{B_{j}}$ is increasing. For each $1\leq i\leq\overline{r}$, compute (by [AM23, Proposition 6.1]) the integer $\overline{k_{i,t}}$ and representation $\overline{\pi_{i,t}}$ for $t$ in the segment $[\overline{B_{i}}+1,-A-1]$ recursively by
    $\begin{cases}\overline{\pi_{i,\overline{B_{i}}+1}}=\pi_{A},\\ \overline{\pi_{i,t}}=D_{\rho|\cdot|^{t}}^{(\overline{k_{i,t}})}(\overline{\pi_{i,t+1}})\text{ is the highest derivative.}\end{cases}$
    If $\overline{k_{i,-A-1}}\neq 0$, then $\pi$ is not of Arthur type and the procedure ends. Set $\overline{k_{i,t}}:=0$ if $t\not\in[\overline{B_{i}}+1,-A-1]$.
    Denote $\overline{K_{i,t}}:=\overline{k_{i,t}}-\overline{k_{i-1,t}}$. For $t\in[\overline{B_{i}}-1,-A-1]$, if $\overline{K_{i,t}}>\overline{K_{i,t+1}}$, then $\pi$ is not of Arthur type and the procedure ends. If $\overline{K_{i,t}}<\overline{K_{i,t+1}}$, then add $\overline{K_{i,t+1}}-\overline{K_{i,t}}$ copies of $\rho\otimes S_{-(t+1)+\overline{B_{i}}+1}\otimes S_{-(t+1)-\overline{B_{i}}+1}$ to $\psi$, and add the same copies of elements in the segment $[-t-1,\overline{B_{i}}]_{\rho}$ in $\Omega^{-}$.
  5. Step 4:
    Let $\Omega=\Omega^{+}+\Omega^{-}$. Denote the multiplicity of $\rho|\cdot|^{t}$ in $\Omega\setminus\Omega(\pi)$ (resp. $\Omega(\pi)\setminus\Omega$) by $m_{1,t}$ (resp. $m_{2,t}$), and let $M_{t}=(m_{1,-t-1}-m_{1,t})+m_{2,t}$.
    For any $t\in[A+1,\epsilon+1]$, if $M_{t}>M_{t-1}$, then $\pi$ is not of Arthur type and the procedure ends. If $M_{t}<M_{t-1}$, add $M_{t}-M_{t-1}$ copies of $\rho\otimes S_{(t-1)+\epsilon+1}\otimes S_{(t-1)-\epsilon+1}$ to $\psi$.
  6. Step 5:
    Construct the local Arthur packet $\Pi_{\psi}$. If there exists $\mathcal{E}$ in this packet such that $\pi(\mathcal{E})=\pi$, then $\pi$ is of Arthur type. Otherwise, $\pi$ is not of Arthur type.
4. The enhanced Shahidi conjecture
In this section, prove the enhanced Shahidi conjecture for symplectic and split odd special orthogonal groups $G_{n}$ (see Theorems 4.7, 4.9).
4.1. A criterion for $|\Psi(\pi)|=1$ for tempered representation
In this subsection, we give a criterion for $|\Psi(\pi)|=1$ for any tempered representation $\pi=\pi(\phi,\varepsilon)$.
First, we state several immediate consequences of the definition of the operators $ui_{i,j},dual_{k}^{+},dual_{k}^{-}$ (see [HLL22, Definitions 3.23, 5.1, 6.5]). Here we recall that two segments $\Delta_{1},\Delta_{2}$ are not linked if $\Delta_{1}\cup\Delta_{2}$ is not a segment or either $\Delta_{1}\supseteq\Delta_{2}$ or $\Delta_{1}\subseteq\Delta_{2}$.
Lemma 4.1.
Let $\mathcal{E}=\cup_{\rho}\{([A_{i},B_{i}]_{\rho},l_{i},\eta_{i})\}_{i\in(I_{\rho},>)}\in\underline{\mathrm{Rep}}$. Let $i,j\in I_{\rho}$ and $i+1>i$ be adjacent.
  1. (a)
    If $[A_{i},B_{i}]_{\rho}$ and $[A_{j},B_{j}]_{\rho}$ are not linked, then $ui_{i,j}$ is not applicable on $\mathcal{E}$.
  2. (b)
    Suppose $A_{i}=B_{i}$ and $A_{i+1}=B_{i+1}$. Then, the operator $ui_{i,i+1}$ is applicable on $\mathcal{E}$ if and only if $A_{i+1}=A_{i}+1$ and $\eta_{i}=-\eta_{i+1}$.
  3. (c)
    Suppose $i$ is minimal with respect to $>$ and $B_{i}=1/2$. Then, $dual_{i}^{+}$ is applicable on $\mathcal{E}$ if and only if $\eta_{i}=-1$.
  4. (d)
    The operator $dual_{i}^{-}$ is applicable on $\mathcal{E}$ only if $B_{i}=-1/2$.
Proof.
All of these follow directly from the definitions of these operators, which we omit the details. ∎
Now, we describe when a tempered representation of good parity lies in exactly one local Arthur packet. Note that these conditions are exactly the opposite of the conditions for $\pi(\phi,\varepsilon)$ being supercuspidal (see [Mœ11a, Theorem 2.5.1] or [Xu17a, Theorem 3.3]).
Theorem 4.2.
Let $\pi=\pi(\phi,\varepsilon)$ be a tempered representation of $G_{n}$ of good parity. Then $\Psi(\pi):=\{\psi\ |\ \pi\in\Pi_{\psi}\}$ is a singleton if and only if the following condition holds.
  1. $\bullet$
    If $\rho\otimes S_{a}\subset\phi$ and $\rho\otimes S_{a+2}\subset\phi$, then $\varepsilon(\rho\otimes S_{a})\varepsilon(\rho\otimes S_{a+2})=1$.
  2. $\bullet$
    If $\rho\otimes S_{2}\subset\phi$, then $\varepsilon(\rho\otimes S_{2})=1$.
Proof.
Write
$\phi=\bigoplus_{\rho}\bigoplus_{i\in I_{\rho}}\rho\otimes S_{2z_{i}+1},$
and let $\mathcal{E}=\cup_{\rho}\{([z_{i},z_{i}]_{\rho},0,\varepsilon(\rho\otimes S_{2z_{i}+1}))\}_{i\in(I_{\rho,>})}$ so that $\pi(\mathcal{E})=\pi$. Here the total order $>$ is non-decreasing with respect to $z_{i}$.
First, we claim that if any of the conditions fails, then there is an operator $T$ applicable on $\mathcal{E}$, and hence $\Psi(\pi)\supseteq\{\psi_{\mathcal{E}}\neq\psi_{T(\mathcal{E})}\}$ is not a singleton by Theorem 3.6(1). Indeed, if the first condition fails, say $z_{i+1}=z_{i}+1$ and $\varepsilon(\rho\otimes S_{2z_{i}+1})=-\varepsilon(\rho\otimes S_{2z_{i}+3})$, then $T=ui_{i,i+1}$ is applicable on $\mathcal{E}$ by Lemma 4.1(b). If the second condition fails, let $i$ be the minimal element in $(I_{\rho},>)$. Then we must have $A_{i}=B_{i}=1/2$ and $\eta_{i}=-1$. Thus, $T=dual_{i}^{+}$ is applicable on $\mathcal{E}$ by Lemma 4.1(c). This completes the verification of the claim.
Next, we show that if the conditions hold, then there is no operator applicable on $\mathcal{E}$, and hence $\Psi(\pi)$ is a singleton by Theorem 3.6(2). We verify that $ui_{i,j},dual_{i}^{+},dual\circ ui_{j,i}\circ dual,dual_{i}^{-}$ are not applicable case by case for any $i<j\in I_{\rho}$.
Suppose the contrary that $ui_{i,j}$ is applicable on $\mathcal{E}$. Lemma 4.1(a) implies that $[z_{i},z_{i}]_{\rho}$ and $[z_{j},z_{j}]_{\rho}$ are linked, which holds only if $z_{j}=z_{i}+1$. However, the first condition in the statement contradicts to the applicability of $ui_{i,j}$ by Lemma 4.1(b). Similarly, suppose the contrary that $dual_{i}^{+}$ is applicable on $\mathcal{E}$. Then $z_{i}=1/2$ and hence we can assume $i$ is minimal in $(I_{\rho},>)$. Again, the second condition in the statement contradicts to the applicability of $dual_{i}^{+}$ by Lemma 4.1(c).
To show $dual\circ ui_{j,i}\circ dual$ is never applicable on $\mathcal{E}$, observe that the corresponding segments in $dual(\mathcal{E})$ are $[z_{i},-z_{i}]$ and $[z_{j},-z_{j}]$, which are never linked. Thus, $ui_{j,i}$ is never applicable on $dual(\mathcal{E})$ by Lemma 4.1(a), and hence $dual\circ ui_{j,i}\circ dual$ is never applicable on $\mathcal{E}$. Finally, none of the $z_{i}$ is equal to $-1/2$, and hence $dual_{i}^{-}$ is never applicable on $\mathcal{E}$ by Lemma 4.1(d). This completes the proof of the theorem. ∎
4.2. The enhanced Shahidi conjecture
In this subsection, we recall and prove the enhanced Shahidi conjecture. Note that we have implicitly fixed a choice of Whittaker datum. That is, we fixed a $G_{n}$-conjugacy class of a tuple $(B,\chi)$, where $B$ is an $F$-rational Borel subgroup of $\mathrm{G}_{n}$ and $\chi$ is a generic character of the $F$-points of the unipotent radical of $B$. For a tempered local Arthur parameter $\psi$, this choice fixes the bijection between $\Pi_{\psi}$ and $\widehat{\mathcal{S}}_{\psi}$ in Theorem 2.11
In [Sha90, Conjecture 9.4], Shahidi conjectured that for any Whittaker data $(B^{\prime},\chi^{\prime}),$ a tempered $L$-packet contains a $(B^{\prime},\chi^{\prime})$-generic representation. For Archimedean fields, Shelstad established a stronger form of this conjecture ([She08]). Namely, for any Whittaker data $(B,\chi),$ a tempered $L$-packet contains a unique $(B,\chi)$-generic representation. Moreover, if the bijection between $\Pi_{\psi}$ and $\widehat{\mathcal{S}}_{\psi}$ is fixed by $(B,\chi),$ then this representation corresponds to the trivial character in $\widehat{\mathcal{S}}_{\psi}$. The non-Archimedean analogue of Shelstad’s result is also true. The Shahidi conjecture is known for symplectic and split odd orthogonal groups $G_{n}$ as follows (see [Kon02, JS03, Liu11, JS12, MW12, Wal12, Art13, Var17]).
Theorem 4.3.
Suppose $\phi$ is a tempered L-parameter of $G_{n}$, then $\pi(\phi,\varepsilon)$ is generic with respect to the Whittaker datum $(B,\chi)$ if and only if the character $\varepsilon$ is trivial.
We recall the enhanced version of the Shahidi conjecture for $\Psi(G_{n})$ now.
Conjecture 4.4 ([LS22, Conjecture 1.5]).
Let $G$ be any quasi-split reductive group defined over a non-Archimedean local field. Assume that there is a theory of local Arthur packets for $G$ as conjectured in [Art89, §6]. For any local Arthur parameter $\psi\in\Psi(G)$, the local Arthur packet $\Pi_{\psi}$ is tempered if and only if it has a generic member.
The enhanced Shahidi conjecture has global applications. If an automorphic representation $\pi$ lies in a global Arthur packet and $\pi_{v}$ is generic for some finite place $v$, then we can obtain information about the global Arthur parameter. More details will be given in §6 (see Theorem 6.2, Conjecture 6.12, and further Theorems 6.13 and 6.15). We remark that certain cases of the global application are already elaborated in the literature. To be explicit, Magaard and Savin ([MS20, Proposition 8.2], $\mathrm{G}=\mathrm{Sp}_{2n}$) and Chen ([Che23, Lemma 4.5], $\mathrm{G}=\mathrm{Sp}_{2n}$, $\mathrm{SO}_{2n+1}$) proved that if a local component of a cuspidal automorphic representation is a Steinberg representation, then its functorial lifting to $\mathrm{GL}$ is a cuspidal representation. Indeed, the key point is that Steinberg representations only live in one local Arthur packet, whose parameter is tempered and irreducible. See [MS20, Proposition 8.2] and Remark 4.11 for more details.
Conjecture 4.4 requires that we prove it for any representation that is generic with respect to some Whittaker datum; however, it is enough to prove it for a fixed Whittaker datum. Indeed, for a tempered local Arthur parameter $\psi,$ the bijection between $\Pi_{\psi}$ and $\widehat{\mathcal{S}}_{\psi}$ in Theorem 2.11, relies on the choice of Whittaker datum. If we let $(B^{\prime},\chi^{\prime})$ be another choice of Whittaker datum, then the bijection between the tempered local Arthur packet $\Pi_{\psi}$ and $\widehat{\mathcal{S}}_{\psi}$ may change; however, Theorem 4.3 shows that $\Pi_{\psi}$ still has a generic member with respect to the Whittaker datum $(B^{\prime},\chi^{\prime}).$
Here is the strategy of the proof of the enhanced Shahidi conjecture for $G_{n}.$ First, irreducible generic representations have been classified for $G_{n}$ (Theorem 4.5). If $\pi$ is of good parity and generic, this allows us to show that $\pi$ must be tempered (Theorem 4.7(i)). By Theorem 4.3, we obtain an extended multi-segment $\mathcal{E}$ for which $\pi(\mathcal{E})=\pi$ and $\psi_{\mathcal{E}}$ is tempered. We check that $\mathcal{E}$ satisfies the conditions of Theorem 4.2 and hence $\pi$ only lies in the tempered local Arthur packet $\Pi_{\psi_{\mathcal{E}}}$.
We proceed with the details now. The following is the classification of generic representations of $G_{n}$ in terms of their $L$-data.
Theorem 4.5 ([Mui98, §4], [JS04, Definition 5.1], [Liu11, Definition 4.14]).
Any irreducible generic representation $\sigma$ of $\mathrm{SO}_{2n+1}(F)$ (resp. $\mathrm{Sp}_{2n}(F)$) is an irreducible parabolic induction of the form (this gives the $L$-data of $\sigma$)
$\sigma=\Delta_{\rho_{1}}[x_{1},-y_{1}]\times\cdots\times\Delta_{\rho_{f}}[x_{f},-y_{f}]\rtimes\pi(\phi,1),$
where the sequence of segments $\{\Sigma_{i}:=[y_{i},-x_{i}]_{\rho_{i}}\}_{i=1}^{f}$ is $\mathrm{SO}_{2n+1}$-generic (resp. $\mathrm{Sp}_{2n}$-generic) with respect to the tempered generic representation $\pi(\phi,1)$.
We do not recall the explicit description of $\mathrm{SO}_{2n+1}$-generic (resp. $\mathrm{Sp}_{2n}$-generic) here. We shall only use the following special case of the definition. Suppose $\{\Sigma_{i}\}_{i=1}^{f}$ is $\mathrm{SO}_{2n+1}$-generic (resp. $\mathrm{Sp}_{2n}$-generic). Then, if $\rho_{i}$ is self-dual of orthogonal type (resp. symplectic type), then $\Sigma_{i}\neq[1/2,1/2]_{\rho_{i}}$. In particular, if $\sigma$ is generic and lies in the local Arthur packet $\Pi_{\psi}$ where $\psi$ is of good parity, then there is no segment of the form $\Delta_{\rho}[-1/2,-1/2]$ in the $L$-data of $\sigma$.
Here are two simple observations.
Lemma 4.6.
  1. (a)
    Suppose $\phi=\sum_{i=1}^{n}\rho_{i}\otimes S_{2z_{i}+1}$ is a tempered L-parameter of $G_{n}$ of good parity. For any $z>1/2$, the order of the highest derivative of $\pi(\phi,1)$ with respect to $D_{\rho|\cdot|^{z}}$ is given by
    $k=\#\{i\ |\ \rho_{i}\cong\rho,z_{i}=z\}$
    and $D_{\rho|\cdot|^{z}}^{(k)}(\pi(\phi,1))=\pi(\phi^{\prime},1)$, where
    $\phi^{\prime}=\phi-(\rho\otimes S_{2z+1})^{\oplus k}+(\rho\otimes S_{2(z-1)+1})^{\oplus k}.$
  2. (b)
    Suppose $\mathcal{E}\in\underline{\mathrm{Rep}}$ and $\mathcal{E}_{\rho}$ is of the form
    $\mathcal{E}_{\rho}=\{([A_{k},B_{k}]_{\rho},l_{k},\eta_{k})\}_{k=1}^{n-r}+\{([y,y]_{\rho},0,\eta_{n})^{r}\},$
    where $A_{k}<y$ for all $1\leq k\leq n-r$. If
    $\pi(\mathcal{E})=L(\Delta_{\rho_{1}}[x_{1},-y_{1}],\dots,\Delta_{\rho_{f}}[x_{f},-y_{f}];\pi(\phi,\varepsilon)),$
    then $x_{i},y_{i}<y$ for $\rho_{i}\cong\rho$, and $\phi$ contains $r$ copies of $\rho\otimes S_{2y+1}$.
Proof.
Part (a) follows directly from [AM23, Theorem 7.1]. For Part (b), it is not hard to see from the definition of $\pi(\mathcal{E})$ that the conclusion holds when $y$ is sufficiently large. The general case follows from taking derivatives (in the sense of [HLL22, Lemma 4.3(iv)]) which changes $\phi$ to $\phi-(\rho\otimes S_{2y+1})^{\oplus r}+(\rho\otimes S_{2(y-1)+1})^{\oplus r}$ at each stage. This completes the proof of the lemma. ∎
Now we are ready to prove Conjecture 4.4.
Theorem 4.7.
Let $G_{n}$ be a symplectic or split odd special orthogonal group.
  1. (i)
    Suppose $\sigma\in\Pi_{\psi}$ for some $\psi\in\Psi(G_{n})$. Then $\sigma$ is generic only if it is tempered.
  2. (ii)
    Any tempered generic representation lives in exactly one local Arthur packet.
Hence, combining with Theorem 4.3, Conjecture 4.4 holds for $G_{n}$.
Proof.
For Part (i), we first deal with the case that $\psi$ is of good parity. In this case, we write $\sigma=\pi(\mathcal{E})$ for $\mathcal{E}=\cup_{\rho}\mathcal{E}_{\rho}\in\underline{\mathrm{Rep}}$.
Suppose for contradiction that $\sigma$ is not tempered. By Theorem 4.5, we write
(4.1) $\sigma=\Delta_{\rho_{1}}[x_{1},-y_{1}]\times\cdots\times\Delta_{\rho_{f}}[x_{f},-y_{f}]\rtimes\pi(\phi,1),$
and
$\phi=\sum_{j=f+1}^{m}\rho_{j}\otimes S_{2z_{j}+1}.$
Let $\rho:=\rho_{1}$ and denote
$y$ $=\max\{y_{i}\ |\ \rho_{i}\cong\rho\},$
$z$ $=\max\left(\{z_{j}\ |\ \rho_{j}\cong\rho\}\cup\{0\}\right).$
Note that for any $1\leq i\leq m$, $y_{i}>x_{i}\geq-y_{i}$ and Theorem 4.5 implies $y_{i}>1/2$. Also, $y\neq 0$. Now we consider two cases.
Case 1. Suppose $z\leq y$. Write $\mathcal{E}_{\rho}=\{([A_{k},B_{k}]_{\rho},l_{k},\eta_{k})\}_{k=1}^{n}$. In this case, we have
$y=\max\{\alpha\ |\ \rho|\cdot|^{\alpha}\in\Omega(\pi(\mathcal{E}))_{\rho}\}.$
Therefore, [HLL22, Theorem 4.10] implies
#$\#\{i\ |\ \rho_{i}\cong\rho,\ y_{i}=y\}+\#\{j\ |\ \rho_{j}\cong\rho,\ z_{j}=y\}=\#\{k\ |\ A_{k}=y\}=\#\{k\ |\ A_{k}\geq y\}.$
On the other hand, from the form of (4.1), the quantity
#$\#\{i\ |\ \rho_{i}\cong\rho,\ y_{i}=y\}+\#\{j\ |\ \rho_{j}\cong\rho,\ z_{j}=y\}$
is precisely the order of the highest derivative of $\sigma$ with respect to $D_{\rho|\cdot|^{y}}$ by Lemma 2.3 and Lemma 4.6(a). Therefore, [Xu17b, Proposition 8.3] implies
#$\#\{i\ |\ \rho_{i}\cong\rho,\ y_{i}=y\}+\#\{j\ |\ \rho_{j}\cong\rho,\ z_{j}=y\}\leq\#\{k\ |\ B_{k}=y\}.$
Then the inequality (from the condition $B_{k}\leq A_{k}$)
#$\#\{k\ |\ A_{k}\geq y\}\geq\#\{k\ |\ B_{k}=y\}$
implies that $\mathcal{E}_{\rho}$ is of the form
$\mathcal{E}_{\rho}=\{([A_{k},B_{k}]_{\rho},l_{k},\eta_{k})\}_{k=1}^{n-r}+\{([y,y]_{\rho},0,1)^{r}\}$
with $A_{k}<y$ for $1\leq k\leq n-r$. However, this contradicts to Lemma 4.6(b).
Case 2. Suppose $z>y$. In this case, we construct a sequence of irreducible generic representations $\{\sigma_{z},\dots,\sigma_{y+1},\sigma_{y}\}$ inductively. For $z\geq x\geq y$ such that $y-x\in\mathbb{Z}$, we define tempered local Arthur parameters $\phi_{i}$ inductively by
  1. $\bullet$
    $\phi_{z}=\phi$,
  2. $\bullet$
    $\pi(\phi_{x-1},1)$ is the highest derivative of $\pi(\phi_{x},1)$ with respect to $\rho|\cdot|^{x}$.
Note that by Lemma 4.6(a), we have
$\max\{\alpha\ |\ \rho\otimes S_{2\alpha+1}\subset\phi_{x}\}=x.$
Now we consider
$\sigma_{x}:=\Delta_{\rho_{1}}[x_{1},-y_{1}]\times\cdots\times\Delta_{\rho_{f}}[x_{f},-y_{f}]\rtimes\pi(\phi_{x},1).$
We claim that if $\sigma_{x}$ is irreducible and generic then so is $\sigma_{x-1}$. Indeed, we have $D_{\rho|\cdot|^{x}}^{(k_{x})}(\pi(\phi_{x},1))=\pi(\phi_{x-1},1)$, so it follows from Lemma 2.3 that $D_{\rho|\cdot|^{x}}^{(k_{x})}(\sigma_{x})=\sigma_{x-1}$, and it is the highest derivative. In particular, $\sigma_{x-1}$ is an irreducible parabolic induction of a product of generic representations, and hence $\sigma_{x-1}$ is also generic, which verifies the claim. Now $\sigma_{y}$ is a generic irreducible representation in Case 1, which gives a contradiction. This completes the proof of Part (i) in the good parity case.
Next, we deal with the general case. By Theorem 2.7, we decompose the local Arthur parameter $\psi=\psi_{np}\oplus\psi_{gp}\oplus\psi_{np}^{\vee}$, where $\psi_{gp}$ is of good parity. Then
$\Pi_{\psi}=\{\tau_{\psi_{np}}\rtimes\pi_{gp}\ |\ \pi_{gp}\in\Pi_{\psi_{gp}}\}$
where if $\psi_{np}=\oplus_{\rho}(\oplus_{i\in I_{\rho}}\rho\otimes S_{a_{i}}\otimes S_{b_{i}})$,
$\tau_{\psi_{np}}=\bigtimes_{\rho}\bigtimes_{i\in I_{\rho}}\begin{pmatrix}\frac{a_{i}-b_{i}}{2}&\cdots&\frac{a_{i}+b_{i}}{2}-1\\ \vdots&\ddots&\vdots\\ -\frac{a_{i}+b_{i}}{2}+1&\cdots&-\frac{a_{i}-b_{i}}{2}\end{pmatrix}_{\rho}.$
Therefore, we may write our generic representation $\pi=\tau_{\psi_{np}}\rtimes\pi_{gp}$.
We know if $\tau\rtimes\sigma$ is irreducible and generic, then $\tau$ and $\sigma$ are necessarily generic. As a consequence, any generalized Speh representation in the product of $\tau_{\psi_{np}}$ and $\pi_{gp}$ is generic. Our argument in the good parity case then implies $\psi_{gp}$ is tempered.
On the other hand, by the classification of the generic dual of $\mathrm{GL}_{d}(F)$, each generalized Speh representation in the product of $\tau_{\psi_{np}}$ is generic if and only if all the segments in its Langlands data are not linked, i.e., it contains only one column. In other words, for any $i$, we have
$\frac{a_{i}-b_{i}}{2}=\frac{a_{i}+b_{i}}{2}-1,$
which implies $b_{i}=1$, and hence $\psi_{np}$ is also tempered. This completes the proof of Part (i).
Part (ii) follows directly from Theorem 4.2. This completes the proof of the theorem and the proof of the enhanced Shahidi conjecture. ∎
Finally, we discuss the case of general local Arthur parameters by extending Conjecture 4.4 to $\Psi^{+}(G)$.
Conjecture 4.8 (Enhanced Shahidi Conjecture for $\Psi^{+}(G)$).
For any quasi-split reductive group $G$, assume that there is a local Arthur packets theory for $G$ as conjectured in [Art89, §6]. For any local Arthur parameter $\psi\in\Psi^{+}(G)$, the local Arthur packet $\Pi_{\psi}$ is generic if and only if it has a generic member.
Suppose $\psi\in\Psi^{+}(G_{n})$ has a decomposition
(4.2) $\psi=\bigoplus_{\rho}\bigoplus_{i\in I_{\rho}}\rho|\cdot|^{x_{i}}\otimes S_{a_{i}}\otimes S_{b_{i}}.$
Recall that we say $\psi$ is generic if, when decomposed as in (4.2), every $b_{i}=1.$ Note that if $\psi\in\Psi(G_{n})$, then $\psi$ is generic if and only if it is tempered. We confirm Conjecture 4.8 in the case of $G_{n}$ as follows.
Theorem 4.9.
A generic representation $\pi$ lies in $\Pi_{\psi}$ for some $\psi\in\Psi^{+}(G_{n})$ if and only if $\psi$ is generic.
Proof.
The proof is the same as the reduction from the general case to the good parity case in the proof of Theorem 4.7 Part (i), which we omit. ∎
By Theorem 4.7 and Corollary 2.10, we obtain the following theorem.
Theorem 4.10.
Suppose that $\pi$ is generic and $\pi\in\Pi_{\psi}$ for some $\psi\in\Psi^{+}(G_{n}).$ Then $\Psi(\pi)=\{\psi\}.$
We end this section with the following remark on Steinberg representations of $G_{n}$ and the proof of the enhanced Shahidi conjecture for Steinberg representations given in [MS20, Proposition 8.2].
Remark 4.11.
  1. (1)
    Let $G$ be a quasi-split group over $F$ with a Borel subgroup $B$. Denote $\mathbf{1}_{P}$ the trivial representation of $P$. The Steinberg representation $\mathrm{St}_{G}$ of $G$ is the unique irreducible subquotient of $\mathrm{ind}_{B}^{G}\mathbf{1}_{B}$ (non-normalized induction) that is not a subquotient of $\mathrm{ind}_{P}^{G}\mathbf{1}_{P}$ for any other parabolic subgroup $P$ of $G$. It is known that $\mathrm{St}_{G}$ is square-integrable, generic, and its Aubert-Zelevinsky dual is the trivial representation.
    For $G_{n}$, we set $N=2n$ if $G_{n}=\mathrm{SO}_{2n+1}(F)$ and $N=2n+1$ if $G_{n}=\mathrm{Sp}_{2n}(F)$. Let $\rho$ be the trivial representation of $W_{F}$ and
    $\phi_{\mathrm{St}}:=\rho\otimes S_{N}.$
    Then the Steinberg representation $\mathrm{St}_{G_{n}}$ is isomorphic to $\pi(\phi_{\mathrm{St}},1)$, the unique representation in the tempered $L$-packet $\Pi_{\phi_{\mathrm{St}}}$. Indeed, there is an injection
    $\pi(\phi_{\mathrm{St}},1)\hookrightarrow\rho|\cdot|^{\frac{N-1}{2}}\times\rho|\cdot|^{\frac{N-3}{2}}\times\cdots\times\rho|\cdot|^{\frac{N+1}{2}-n}\rtimes\mathbf{1}_{G_{0}},$
    which shows that $D(\pi(\phi_{\mathrm{St}},1))\neq 0$ (see [HLL22, Lemma 2.5]), where
    $D:=D_{\rho|\cdot|^{\frac{N-1}{2},\dots,\frac{N+1}{2}-n}},$
    and $\pi(\phi_{\mathrm{St}},1)$ is a subquotient of $\mathrm{ind}_{B}^{G_{n}}\mathbf{1}_{B}$. On the other hand, $D(\mathrm{ind}_{P}^{G_{n}}\mathbf{1}_{P})=0$ for any other parabolic subgroup $P$ of $G_{n}$ by Lemma 2.3 and the fact that
    $D_{\rho|\cdot|^{x}}(\mathbf{1}_{G_{m}})=0$
    for any $x>0$ and $m\in\mathbb{Z}_{>0}$. Therefore, $\pi(\phi_{\mathrm{St}},1)=\mathrm{St}_{G_{n}}$.
  2. (2)
    The enhanced Shahidi conjecture implies $\Psi(\mathrm{St}_{G_{n}})$ is a singleton. This is already proven by Magaard and Savin (((([MS20, Proposition 8.2])))) using a different approach based on the following property of $\phi_{\mathrm{St}}$. Let $\psi_{\mathrm{St}}=\phi_{St}\otimes S_{1}$. We have
    $\{\psi\ |\ \psi^{\Delta}=(\psi_{\mathrm{St}})^{\Delta}\}=\{\psi_{\mathrm{St}},\widehat{\psi_{\mathrm{St}}}\},$
    where $\psi^{\Delta}$ is an $L$-parameter defined by $\psi^{\Delta}(w,x):=\psi(w,x,x)$. Thus [Mœ09a, Corollary 4.2] implies $\Psi(\mathrm{St}_{G_{n}})\subseteq\{\psi_{\mathrm{St}},\widehat{\psi_{\mathrm{St}}}\}$. However, since the trivial representation $\mathbf{1}_{G_{n}}$ is not tempered, it is not in $\Pi_{\psi_{\mathrm{St}}}$, and hence $\mathrm{St}_{G_{n}}=\widehat{\mathbf{1}_{G_{n}}}$ is not in $\Pi_{\widehat{\psi_{\mathrm{St}}}}$. Therefore, $\Psi(\mathrm{St}_{G_{n}})=\{\psi_{\mathrm{St}}\}$.
5. Unramified representations of Arthur type
In this section, we classify representations of Arthur type of $G_{n}$ which are unramified with respect to $\mathrm{G}_{n}(\mathcal{O}_{F})$ (see Theorem 5.2) and consider its applications. These unramified representations are important for global considerations, e.g., the applications in §6, and hence it is desirable to determine all the local Arthur packets to which an unramified representation of Arthur type belongs. Towards this goal, we make the following conjecture, predicting that these unramified representations live in only one local Arthur packet, which is anti-generic, i.e., the parameter is trivial on the Deligne-$\mathrm{SL}_{2}(\mathbb{C})$. We confirm the conjecture for $G_{n}$ in Corollary 5.3.
Conjecture 5.1.
Let $G$ be a connected reductive group defined over a non-Archimedean local field. Assume that there is a local Arthur packets theory for $G$ as conjectured in [Art89, §6]. Then any unramified representation of $G$ of Arthur type lies in exactly one local Arthur packet. Moreover, it lies in the $L$-packet associated to an anti-generic local Arthur packet.
Now we start the classification of all unramified representations of Arthur type. Let $\pi$ be a representation of Arthur type and write $\pi=\tau_{nu,>0}\times\tau_{np}\rtimes\pi_{gp}$ by Theorem 2.8. Then, $\pi$ is unramified if and only if each of $\tau_{nu,>0},\tau_{np}$ and $\pi_{gp}$ is unramified. Thus, the problem is reduced to the classification of unramified representations of Arthur type and of good parity. We answer this question in the following theorem, which will be proved by applying Algorithm 3.9. Note that if $\pi\in\Pi_{\psi}$ for some $\psi\in\Psi^{+}(G_{n})$ and $\pi$ is of good parity, then $\psi\in\Psi(G_{n})$ automatically by Theorem 2.8 and the definition of good parity (Definition 2.9).
Theorem 5.2.
Suppose $\pi$ is an irreducible representation of $G_{n}$ of good parity, written as
$\pi=L(\Delta_{\rho_{1}}[x_{1},y_{1}],\dots,\Delta_{\rho_{f}}[x_{f},y_{f}];\pi(\phi,\varepsilon)),$
where
$\phi=\sum_{i=f+1}^{m}\rho_{i}\otimes S_{2z_{i}+1}.$
Then $\pi$ is unramified and of Arthur type if and only if the followings hold.
  1. (i)
    $x_{i}=y_{i}$ for $1\leq i\leq f$, $z_{i}=0$ for $f+1\leq i\leq m$, and $\rho_{i}$ are unramified characters for $1\leq i\leq m$.
  2. (ii)
    For any $\rho$ and $x\geq 0$, define
    $m_{\rho,x}:=\begin{cases}\#\{1\leq i\leq f\ |\ \rho_{i}\cong\rho,x_{i}=-x\}&\text{if }x>0,\\ \#\{f+1\leq i\leq m\ |\ \rho_{i}\cong\rho\}&\text{if }x=0.\end{cases}$
    Then for any $\rho$ and $x\geq 0$, $m_{\rho,x+1}\leq m_{\rho,x}$.
  3. (iii)
    The character $\varepsilon$ is trivial.
Moreover, in this case, we have $\pi=\pi(\mathcal{E})$ where (set $m_{\rho,x}=0$ for $x<0$)
(5.1) $\mathcal{E}=\cup_{\rho}\{([x,-x]_{\rho},\lfloor x+1/2\rfloor,1)^{m_{\rho,x-1}-m_{\rho,x}}\}_{x\in\frac{1}{2}\mathbb{Z}}.$
Here the exponent $m_{\rho,x-1}-m_{\rho,x}$ denotes the multiplicity of the extended segment.
Proof.
Suppose $\pi$ satisfies all the conditions. Then it follows from [HLL22, Theorem 9.5] that $\pi(\mathcal{E})=\pi$, and hence $\pi$ is of Arthur type. Condition (i) implies that the $L$-parameter $\phi_{\pi}=\phi_{\psi_{\mathcal{E}}}$ is unramified. Thus the $L$-packet $\Pi_{\phi_{\psi_{\mathcal{E}}}}$ contains an unramified representation by the Satake isomorphism ([Sat63]). According to [Art13, Theorem 1.5.1(a)], this unramified representation in the local Arthur packet $\Pi_{\psi_{\mathcal{E}}}$ must correspond to the trivial character in $\widehat{\mathcal{S}}_{\psi}$. On the other hand, [Ato22b, Theorem 3.6] confirms that $\pi(\mathcal{E})$ is the only representation in $\Pi_{\psi}$ corresponding to the trivial character in $\widehat{\mathcal{S}}_{\psi}$. Therefore, we conclude that this unramified representation is exactly $\pi=\pi(\mathcal{E})$.
Next, we show that if $\pi$ is unramified and of Arthur type, then it satisfies Conditions (i), (ii) and (iii). Condition (i) follows from $\phi_{\pi}$ being unramified. Assuming Condition (ii) is already verified, then by [HLL22, Theorem 9.5], we have $\pi=\pi(\mathcal{E}^{\prime})$, where
$\mathcal{E}^{\prime}=\cup_{\rho}\{([x,-x]_{\rho},\lfloor x+1/2\rfloor,\varepsilon_{\rho})^{m_{\rho,x-1}-m_{\rho,x}}\}_{x\in\frac{1}{2}\mathbb{Z}},$
with $\varepsilon_{\rho}:=\varepsilon(\rho\otimes S_{1})$ if $m_{\rho,0}>0$ and $\varepsilon_{\rho}:=1$ otherwise. Again [Ato22b, Theorem 3.6] and [Art13, Theorem 1.5.1(a)] show that $\pi$ is unramified only if $\varepsilon_{\rho}=1$ for all $\rho$, which verifies Condition (iii). Therefore, it remains to show that $\pi$ is of Arthur type only if Condition (ii) holds. We confirm this by applying Algorithm 3.9 on $\pi$.
[…]
For Step 3, it follows from [AM23, Proposition 6.1] that if $\pi$ satisfies Condition (i), then for $0<x\leq A$,
$D_{\rho|\cdot|^{-x}}^{(m_{\rho,x}-m_{\rho,x+1})}(\pi)$
is the highest derivative. Moreover, the resulting representation also satisfies Condition (i). From this observation, we compute the ingredients in Step 3 of the algorithm as follows.
  1. $\bullet$
[…]
  2. $\bullet$
    $\overline{k_{i,t}}=\max(m_{\rho,t}-m_{\rho,t+1},0)$.
  3. $\bullet$
    $\overline{K_{i,t}}=m_{\rho,-\overline{B_{i}}}-m_{\rho,-(\overline{B_{i}}-1)}$ if $t=\overline{B_{i}}$, and $\overline{K_{i,t}}=0$ $t<\overline{B_{i}}$.
Thus at the end of Step 3, we obtain
$\psi=\bigoplus_{0<x\leq A}(\rho\otimes S_{1}\otimes S_{2x+1})^{\max(m_{\rho,x}-m_{\rho,x+1},0)},$
and
$\Omega=\sum_{0<x\leq A}([x,-x]_{\rho})^{\max(m_{\rho,x}-m_{\rho,x+1},0)}.$
Here in the direct sum and summation, $x$ is assumed to be in $\mathbb{Z}+A$.
Note that for $x\in A+\mathbb{Z}\setminus\{0\}$, the multiplicity of $\rho|\cdot|^{x}$ in $\Omega$ is given by
$\sum_{|x|\leq y\leq X}\max(m_{\rho,y}-m_{\rho,y+1},0)\geq m_{\rho,|x|},$
and hence $\Omega(\pi)\setminus\Omega=\{(\rho)^{\max(m_{\rho,0}-m_{\rho,1},0)}\}.$
In Step 4, we suppose the contrary that there exists an $x$ such that $m_{\rho,x}<m_{\rho,x+1}$. Note that this implies $x+2\leq A+1$. Choose a maximal $x$ with this property. Then the multiplicities of $\rho|\cdot|^{\pm(x+2)}$ (resp. $\rho|\cdot|^{\pm(x+1)}$, $\rho|\cdot|^{\pm x}$) in $\Omega$ and $\Omega(\pi)$ are $m_{\rho,x+2}$ (resp. $m_{\rho,x+1}$, $m_{\rho,x+1}$) and $m_{\rho,x+2}$ (resp. $m_{\rho,x+1}$, $m_{\rho,x}$) respectively. In any case, we have
$M_{x+1}=0>-(m_{\rho,x+1}-m_{\rho,x})=M_{x},$
and hence $\pi$ is not of Arthur type. This completes the verification of Condition (ii) and the proof of the theorem. ∎
The above theorem has several applications. First, it implies the Aubert-Zelevinsky dual of these unramified representations are tempered. We remark that this gives another proof of [MT11, Theorem 0-5]. Second, we prove Conjecture 5.1 for $G_{n}$ as follows.
Corollary 5.3.
Conjecture 5.1 holds for $G_{n}.$ That is, if $\pi$ is unramified of Arthur type, then $\Psi(\pi)=\{\psi\}$ is a singleton, $\pi\in\Pi_{\phi_{\psi}}$, and $\psi$ is anti-generic.
Proof.
Suppose $\pi$ is unramified of Arthur type. Then write $\pi=\tau_{nu,>0}\times\tau_{np}\rtimes\pi_{gp}$ as in Theorem 2.8. Since $\phi_{\pi}$ decomposes as a direct sum of one-dimensional representations, so does $\phi_{\tau_{nu,>0}}=\phi_{\psi_{nu,>0}}$ and $\phi_{\tau_{np}}=\phi_{\psi_{np}}$. This shows that the restrictions of $\psi_{nu,>0}$ and $\psi_{np}$ to Deligne-$\mathrm{SL}_{2}(\mathbb{C})$ are trivial. Thus by Corollary 2.10, it suffices to show $\Psi(\pi_{gp})$ is a singleton consisting of an anti-tempered local Arthur parameter.
Now we apply Theorem 5.2 to $\pi_{gp}$. Take $\mathcal{E}$ as in (5.1) with $\pi(\mathcal{E})=\pi_{gp}$. We have
$dual(\mathcal{E})=\cup_{\rho}\{([x,x]_{\rho},0,\varepsilon_{\rho})^{m_{\rho,x-1}-m_{\rho,x}}\}_{x\in\frac{1}{2}\mathbb{Z}},$
where $\varepsilon_{\rho}$ only depends on $\rho$ and $\varepsilon_{\rho}=1$ if $\rho\otimes S_{2}$ is of the same parity as $\widehat{G}_{n}$. Note that
$\psi_{dual(\mathcal{E})}=\bigoplus_{\rho}\bigoplus_{x\in\frac{1}{2}\mathbb{Z}}(\rho\otimes S_{2x+1}\otimes S_{1})^{m_{\rho,x-1}-m_{\rho,x}},$
which is tempered, and hence $\widehat{\pi}_{gp}=\widehat{\pi(\mathcal{E})}=\pi(dual(\mathcal{E}))$ is tempered. Moreover, we have $\widehat{\pi}_{gp}=\pi(\phi_{\psi_{dual(\mathcal{E})}},\varepsilon)$, where $\varepsilon(\rho\otimes S_{2x+1})=\varepsilon_{\rho}$. In particular, $\widehat{\pi}_{gp}$ satisfies all the conditions of Theorem 4.2. Hence $\Psi(\widehat{\pi(\mathcal{E})})=\{\psi_{dual(\mathcal{E})}\}$. This implies $\Psi(\pi_{gp})=\{\psi_{\mathcal{E}}\}$, which completes the proof of the corollary. ∎
Remark 5.4.
The most crucial condition we used for an unramified representation $\pi$ is that $\phi_{\pi}$ decomposes as a direct sum of one-dimensional representations. In fact, if $\pi$ is of Arthur type and of good parity and has this property, then the arguments in Theorem 5.2(ii) also work. We also obtain $\pi=\pi(\mathcal{E})$ with $\mathcal{E}$ of the form (5.1) with possibly some different $\eta_{i}$ since Part (iii) of the proposition may not hold. However, $\pi(\mathcal{E})$ still lies in $\Pi_{\phi_{\psi}}$ for some anti-tempered $\psi$ by [HLL22, Theorem 9.5]. Then the same arguments in Corollary 5.3 work, and we conclude that $\Psi(\pi)=\{\psi\}$. This can be generalized to an arbitrary representation of Arthur type whose $L$-parameter decomposes as a direct sum of one-dimensional representations.
As another application, we prove the following corollary.
Corollary 5.5.
For any $\psi\in\Psi^{+}(G_{n})$, $\Pi_{\psi}$ contains at most one unramified representation.
Proof.
By Theorem 2.8, we may assume $\psi$ is of good parity. Let $\pi_{1},\pi_{2}$ be unramified representations in $\Pi_{\psi}$. Theorem 5.2 gives $\pi_{i}=\pi(\mathcal{E}_{i})$ where $\mathcal{E}_{i}$ is given by (5.1). Then Corollary 5.3 implies that $\psi=\psi_{\mathcal{E}_{1}}=\psi_{\mathcal{E}_{2}}$, and hence $\mathcal{E}_{1}=\mathcal{E}_{2}$. We conclude that $\pi_{1}=\pi_{2}$, which proves the corollary. ∎
We end this subsection by giving some examples of unramified representation of Arthur type.
Example 5.6.
Let $\rho$ be the trivial representation.
  1. 1.
    Consider $\psi=\rho\otimes S_{1}\otimes S_{4}+\rho\otimes S_{1}\otimes S_{2}$, which is a local Arthur parameter of good parity for $\mathrm{SO}_{7}(F).$ We have $\Pi_{\psi}=\{\pi(\mathcal{E}_{1}),\pi(\mathcal{E}_{2})\}$, where
[…]
    We have
    $\pi(\mathcal{E}_{1})$ $=L(\Delta_{\rho}[\frac{-3}{2},\frac{-3}{2}],\Delta_{\rho}[\frac{-1}{2},\frac{-1}{2}],\Delta_{\rho}[\frac{-1}{2},\frac{-1}{2}]),$
    $\pi(\mathcal{E}_{2})$ $=L(\Delta_{\rho}[\frac{-3}{2},\frac{-3}{2}],\Delta_{\rho}[\frac{-1}{2},\frac{-1}{2}];\pi(\frac{1}{2}^{+})).$
    By [HLL22, Theorem 9.5], $\pi(\mathcal{E}_{1})\in\Pi_{\phi_{\psi}}$ and $\pi(\mathcal{E}_{2})\not\in\Pi_{\phi_{\psi}}.$ Note that $\phi_{\psi}$ is an unramified $L$-parameter. Since $\Pi_{\phi_{\psi}}$ is a singleton, by the Satake isomorphism ([Sat63]), $\pi(\mathcal{E}_{1})$ is unramified. We have
[…]
    and $\widehat{\pi(\mathcal{E}_{1})}=\pi(dual(\mathcal{E}_{1}))$. From Theorem 4.3, $\pi(dual(\mathcal{E}_{1}))$ is generic and hence by Theorem 4.7, it lies in exactly one local Arthur packet. Finally, from compatibility of local Arthur packets with the Aubert-Zelevinsky dual ([Xu17b, §A]), we have that the unramified representation $\pi(\mathcal{E}_{1})$ lies in exactly one local Arthur packet.
  2. 2.
    Let $\chi$ be the unique unramified quadratic character of $F$. Consider
    $\psi=\rho\otimes S_{1}\otimes S_{1}+\chi\otimes S_{1}\otimes S_{1}+\chi\otimes S_{1}\otimes S_{1},$
    a local Arthur parameter of $\mathrm{Sp}_{2}(F),$ which is both tempered and anti-tempered. The local Arthur packet $\Pi_{\psi}$ contains two representations $\pi^{+},\pi^{-}$ given by the extended multi-segments
[…]
    Note that $\pi^{+}$ is unramified (with respect to $\mathrm{Sp}_{2}(\mathcal{O}_{F})$) and generic. Since $dual(\mathcal{E}^{+})=\mathcal{E}^{-}$, $\widehat{\pi^{+}}=\pi^{-}$, which is not generic (with respect to the fixed Whittaker datum.) Nevertheless, by Theorem 4.2, $\Psi(\pi^{+})=\{\psi\}=\Psi(\pi^{-})$.
6. Global application
In this section, we discuss the global applications of the results in §4, 5. In particular, we prove Theorems 1.3, 1.5 (Theorems 6.2, 6.13 below), and consider a sequence of Clozel’s conjectures (Conjectures 6.6, 6.10 and 6.14 below).
6.1. Global Arthur packets
In this subsection, we recall the setting of global Arthur packets for $\mathrm{G}_{n}$ ([Art13, §1.5]). Then, as an application of the enhanced Shahidi conjecture (Theorems 4.9, 4.10), we show that for any automorphic representation $\pi\in\Pi_{\psi}$, if $\pi_{v}$ is generic at one finite place, then $\pi_{v}$ is generic for almost all finite places (see Theorem 6.2 below). This result has its own interest and plays a role towards the generalized Ramanujan conjecture of $\mathrm{G}_{n}$ (see Theorem 6.13).
Let $k$ be a number field with ring of adeles $\mathbb{A}_{k}$. Following [Art13, §1.4], a global Arthur parameter $\psi$ is a finite sum of irreducible representations $\mu\otimes S_{b}$ where $\mu$ is an irreducible cuspidal representation of a general linear group over $\mathbb{A}_{k}.$ That is, $\psi=\boxplus_{i=1}^{m}\mu_{i}\otimes S_{b_{i}}$ where $\mu_{i}$ are irreducible cuspidal representations of $\mathrm{GL}_{n_{i}}(\mathbb{A}_{k}).$ Let $v$ be a finite place of $k.$ Then
$(\mu_{i})_{v}=\times_{j\in I_{\mu_{i}}}\mathrm{St}(\tau_{j},a_{j}).$
Here, $I_{\mu_{i}}$ is a finite indexing set, $\tau_{j}$ is an irreducible supercuspidal representation (not necessarily unitary) of a general linear group over $k_{v},$ $a_{i}$ is a positive integer, and $\mathrm{St}(\tau_{i},a_{i})$ is the unique irreducible subrepresentation of
$\tau_{i}|\cdot|^{\frac{a_{i}-1}{2}}\times\cdots\times\tau_{i}|\cdot|^{\frac{1-a_{i}}{2}}.$
We identify $(\mu_{i})_{v}$ with the representation $\oplus_{j\in I_{\mu_{i}}}\tau_{j}\otimes S_{a_{j}}$ of $W_{k_{v}}\times\mathrm{SL}_{2}(\mathbb{C})$. By an approximation of the Ramanujan conjecture, we may assume that $\tau_{j}=\tau_{j}^{\prime}|\cdot|^{x}$ where $\tau_{j}^{\prime}$ is an irreducible unitary supercuspidal representation of a general linear group over $k_{v}$ and $\frac{-1}{2}<x<\frac{1}{2}.$ Thus, for any finite place $v$, we associate to $\psi$ a local Arthur parameter by
(6.1) $\psi_{v}=\bigoplus_{i=1}^{m}\left(\bigoplus_{j\in I_{\mu_{i}}}\tau_{j}\otimes S_{a_{j}}\right)\otimes S_{b_{i}}.$
Say $\psi$ is a global Arthur parameter of $\mathrm{G}_{n}({\mathbb{A}}_{k})$, each local Arthur parameter $\psi_{v}$ factors through $\widehat{\mathrm{G}}_{n}(\mathbb{C})$ and hence $\psi_{v}\in\Psi^{+}(\mathrm{G}_{n}(k_{v}))$.
Arthur constructed global Arthur packets $\Pi_{\psi}$ in [Art13, §1.5]. According to the construction, if $\pi\in\Pi_{\psi}$ with $\pi=\otimes_{v}\pi_{v}$, then $\pi_{v}\in\Pi_{\psi_{v}}$ for all places $v$. Moreover, the character associated to $\pi_{v}$ in $\widehat{\mathcal{S}}_{\psi_{v}}$ is trivial for almost all $v$.
Corollary 6.1.
Suppose that $\pi=\otimes\pi_{v}$ lies in a global Arthur packet $\Pi_{\psi}$ of $\mathrm{G}_{n}$. There exists a finite set of places $\mathcal{S}$, containing the Archimedean places, such that for $v\not\in\mathcal{S},$ $\pi_{v}$ is unramified and $\pi_{v}\in\Pi_{\phi_{\psi_{v}}}.$
Proof.
Let $\mathcal{S}$ be a finite set of places containing the Archimedean places such that for $v\not\in\mathcal{S},$ $(\mu_{i})_{v}$ is unramified for any $1\leq i\leq m$, and the character of $\pi_{v}$ in $\widehat{\mathcal{S}}_{\psi_{v}}$ is trivial. Then for $v\not\in\mathcal{S}$, $\psi_{v}$ is anti-generic and $\phi_{\psi_{v}}$ is unramified. This implies the $L$-packet $\Pi_{\phi_{\psi_{v}}}$ contains an unramified representation $\pi^{ur}_{v}$. By [Art13, Theorem 1.5.1(a)], $\pi^{ur}_{v}\in\Pi_{\psi_{v}}$ corresponds to the trivial character in $\widehat{\mathcal{S}}_{\psi_{v}}$, and hence $\pi^{ur}_{v}=\pi_{v}$ since the map $\Pi_{\psi_{v}}\to\widehat{\mathcal{S}}_{\psi_{v}}$ is a bijection (by construction of $\Pi_{\psi_{v}}$ using extended multi-segments or [Art13, Lemma 7.1.1]). This completes the proof of the corollary. ∎
Now we prove the following application of the enhanced Shahidi conjecture (Theorems 4.9, 4.10).
Theorem 6.2.
Suppose that $\pi=\otimes\pi_{v}$ lies in a global Arthur packet $\Pi_{\psi}$ of $\mathrm{G}_{n}$ and that there exists a finite place $v_{0}$ such that $\pi_{v_{0}}$ is generic. Then the followings hold:
  1. (1)
    The global Arthur parameter is of the form $\psi=\boxplus_{i=1}^{m}\mu_{i}\otimes S_{1},$ where $\mu_{i}$ are irreducible cuspidal representations of general linear groups.
  2. (2)
    For any place $v,$ $\psi_{v}$ is generic. In particular, $\pi_{v}$ is generic for almost all finite places $v$.
  3. (3)
    Furthermore, for almost all places $v$, the local Arthur parameter $\psi_{v}$ has trivial restrictions to both the Deligne-$\mathrm{SL}_{2}(\mathbb{C})$ and Arthur-$\mathrm{SL}_{2}(\mathbb{C}).$
Proof.
Since $\pi_{v_{0}}$ is generic of Arthur type, Theorem 4.10 implies $\Psi(\pi_{v_{0}})$ is a singleton, and hence $\Psi(\pi_{v_{0}})=\{\psi_{v_{0}}\}$. Then, Theorem 4.9 implies $\psi_{v_{0}}$ is generic, and hence it is of the form
$\psi_{v_{0}}=\bigoplus_{\rho}\bigoplus_{i\in I_{\rho}}\rho\otimes S_{a_{i}}\otimes S_{1},$
where $\rho$ is a supercuspidal representation of $\mathrm{GL}(k_{v})$. Comparing with (6.1), we obtain $\psi=\boxplus_{i=1}^{m}\mu_{i}\otimes S_{1}.$ This proves Part (1).
By Part (1), for each finite place $v$, we have
$\psi_{v}=\bigoplus_{i=1}^{m}(\mu_{i})_{v}\otimes S_{1}.$
This proves the first statement in Part (2). Since for almost all finite places, the character of $\pi_{v}$ in $\widehat{\mathcal{S}}_{\psi_{v}}$ is trivial, $\pi_{v}$ is generic by Theorems 4.3, 4.5. This proves the second statement of Part (2). For almost all places $v$, $(\mu_{i})_{v}$ are unramified for all $1\leq i\leq m$, and hence the restriction of $(\mu_{i})_{v}$ to the Deligne-$\mathrm{SL}_{2}(\mathbb{C})$ is trivial for these places. This shows Part (3) and completes the proof of the theorem. ∎
We expect that the result in Theorem 6.2 Part (2) is true for general quasi-split reductive groups, see Conjecture 6.12.
6.2. Clozel’s conjectures
In this subsection, we consider a list of Clozel’s conjectures in [Clo07] on understanding the local components of automorphic representations in the discrete spectrum of reductive groups (see Conjectures 6.6, 6.8, and 6.10 below). As applications of our results in §5 on the characterization of unramified representations of Arthur type, for groups $\mathrm{G}_{n}$, we prove Conjecture 6.8 (see Theorem 6.9 below) and show that the Ramanujan conjecture of $\mathrm{GL}_{n}$ implies Conjectures 6.6 and 6.10 (see Theorems 6.7 and 6.11 below).
Let $k$ be a number field and $\mathrm{G}$ be a reductive group defined over $k$. Let
$\mathcal{A}_{\mathrm{G},disc}=L^{2}_{disc}(\mathrm{G}(k)\backslash\mathrm{G}({\mathbb{A}}_{k}),\omega),$
where $\omega$ is a unitary character of $Z(k)\backslash Z({\mathbb{A}}_{k})$, $Z$ is the center of $\mathrm{G}$. First, we recall Arthur’s classification for the discrete automorphic spectrum for quasi-split classical groups.
Theorem 6.3 ([Art13, Theorem 1.5.2]).
If $\mathrm{G}$ is a symplectic group or quasi-split special orthogonal group, then
$L^{2}_{disc}(\mathrm{G}(k)\backslash\mathrm{G}({\mathbb{A}}_{k}))=\bigoplus_{\psi}\bigoplus_{\pi\in\Pi_{\psi}}m_{\pi}\pi,$
where $\psi$ runs through all global Arthur parameters of $\mathrm{G}({\mathbb{A}}_{k})$, $\Pi_{\psi}$ is the global Arthur packets, and $m_{\pi}$ is the multiplicity.
Next, we recall the Ramanujan conjecture for $\mathrm{GL}_{n}$.
Conjecture 6.4.
If $\pi$ is a cuspidal automorphic representation of $\mathrm{GL}_{n}({\mathbb{A}}_{k})$, then $\pi_{v}$ is tempered for any place $v$.
For each place $v$, recall that $\Psi(\mathrm{G}(k_{v}))$ is the collection of local Arthur parameter of $\mathrm{G}(k_{v})$ whose restriction to Weil group $W_{k_{v}}$ is bounded. We give a direct consequence of Conjecture 6.4.
Lemma 6.5.
Assume Conjecture 6.4, then any local Arthur parameter of $\mathrm{G}(k_{v})$ lies in $\Psi(\mathrm{G}(k_{v}))$.
Proof.
Suppose a local Arthur parameter $\psi_{v}$ comes from the localization of $\psi=\boxplus_{i=1}^{m}\mu_{i}\otimes S_{b_{i}}$. Assuming Conjecture 6.4 holds, then $(\mu_{i})_{v}$ is tempered, and hence its $L$-parameter has bounded image when restricted to $W_{k_{v}}$. Thus the restriction of
$\psi_{v}=\bigoplus_{i=1}^{m}(\mu_{i})_{v}\otimes S_{b_{i}}$
to $W_{k_{v}}$ is also bounded. This completes the proof of the lemma. ∎
Now we are ready to recall Clozel’s conjectures in [Clo07]. The first conjecture describes the Satake parameters of unramified local components of automorphic representations of $\mathrm{G}({\mathbb{A}}_{k})$ in the discrete spectrum.
Conjecture 6.6 ([Clo07, Conjecture 2 (Arthur + $\varepsilon$)]).
For any finite place $v$, if $\pi_{v}$ is an unramified local component of $\pi\in\mathcal{A}_{\mathrm{G},disc}$, then the Satake parameter of $\pi_{v}$ is of the form $\phi_{\psi_{v}}(\mathrm{Frob}_{v})$ for some $\psi_{v}\in\Psi(\mathrm{G}(k_{v}))$. Here $\phi_{\psi_{v}}$ is defined as in (1.1).
As an application of the results in §5, we show that for groups $\mathrm{G}_{n}$, Conjecture 6.6 is implied by the Ramanujan conjecture for $\mathrm{GL}_{n}$.
Theorem 6.7.
When $\mathrm{G}=\mathrm{G}_{n}$, Conjecture 6.4 implies Conjecture 6.6.
Proof.
By Theorem 6.3, $\pi_{v}\in\Pi_{\psi_{v}}$ for some local Arthur parameter $\psi_{v}$. By Corollary 5.3, $\pi_{v}\in\Pi_{\phi_{\psi_{v}}}$, and hence $\phi_{\psi_{v}}(\mathrm{Frob}_{v})$ gives the Satake parameter of $\pi_{v}$. Finally, assuming Conjecture 6.4 holds, Lemma 6.5 implies $\psi_{v}\in\Psi(\mathrm{G}_{n}(k_{v}))$. This completes the proof of the theorem. ∎
We remark that [Clo07, Conjecture 3] follows from [Clo07, Conjecture 2]. Next, we recall [Clo07, Conjecture 2A] which is the reciprocal of Conjecture 5.1, and prove it for the groups $\mathrm{G}_{n}$ without assumption.
Conjecture 6.8 ([Clo07, Conjecture 2A]).
Suppose that $\psi$ is a local Arthur parameter of $G$ such that $\phi_{\psi}$ is unramified. Then the local Arthur packet $\Pi_{\psi}$ contains a unique unramified representation. More specifically, the unramified representation is the one associated to $\phi_{\psi}$ via the Satake isomorphism (((([Sat63])))).
Theorem 6.9.
Conjecture 6.8 holds for groups $\mathrm{G}_{n}$.
Proof.
By Corollary 5.5, $\Pi_{\psi_{v}}$ contains at most one unramified representation, which must be in the unramified $L$-packet $\Pi_{\phi_{\psi_{v}}}$. This completes the proof of the theorem. ∎
The following conjecture of Clozel ([Clo07, Conjecture 4]) can be regarded as a variant of the generalized Ramanujan conjecture for automorphic representations in the discrete spectrum of general reductive groups $\mathrm{G}$, towards the generalized Ramanujan problem (see [Sar05]).
Conjecture 6.10 ([Clo07, Conjecture 4]).
Suppose $\pi=\otimes\pi_{v}\in\mathcal{A}_{\mathrm{G},disc}$.
  1. (1)
    If there exists a finite place ${v_{0}}$ such that $\pi_{v_{0}}$ is unramified and tempered. Then every component of $\pi$ is tempered.
  2. (2)
    In general, for any finite place $v$ such that $\pi_{v}$ is unramified, write the multi-set of the absolute value of eigenvalues of $\phi_{\psi_{v}}(\mathrm{Frob}_{v})$ as $\{q_{v}^{w_{v,1}},\ldots,q_{v}^{w_{v,N}}\}$, where $q_{v}$ is the cardinality of the residue field of $k_{v}$. Then the multi-set $\{w_{v,1},\ldots,w_{v,N}\}$ is independent of the unramified place $v$.
As an application of Corollary 5.3 and Theorem 6.2, we prove that Conjecture 6.4 also implies Conjecture 6.10 for groups $\mathrm{G}_{n}$.
Theorem 6.11.
When $\mathrm{G}=\mathrm{G}_{n}$, Conjecture 6.4 implies Conjecture 6.10.
Proof.
Theorem 6.3 implies $\pi\in\Pi_{\psi}$ for some global Arthur parameter $\psi$. For any finite place $v$ such that $\pi_{v}$ is unramified, Corollary 5.3 implies $\Psi(\pi_{v})=\{\psi_{v}\}$, $\pi_{v}\in\Pi_{\phi_{\psi_{v}}}$, and [Art13, Theorem 1.5.1(a)] implies the character of $\pi_{v}$ in $\widehat{\mathcal{S}}_{\psi_{v}}$ is trivial.
For Part (1), since $\pi_{v_{0}}$ is also tempered, $\Psi(\pi_{v_{0}})$ must contain a tempered member, and hence $\psi_{v_{0}}$ is tempered. As $\pi_{v_{0}}$ corresponds to the trivial character in $\widehat{\mathcal{S}}_{\psi_{v_{0}}}$, Theorem 4.3 implies $\pi_{v_{0}}$ is generic. Now we apply Theorem 6.2(1) to obtain that $\psi=\boxplus_{i=1}^{m}\mu_{i}\otimes S_{1}$, and hence for any place $v$,
$\psi_{v}=\bigoplus_{i=1}^{m}(\mu_{i})_{v}\otimes S_{1}.$
Then, Conjecture 6.4 implies the restriction of $(\mu_{i})_{v}$ to $W_{k_{v}}$ is bounded for any $i$, and hence the restriction $\psi_{v}$ to $W_{k_{v}}$ is also bounded. We conclude that $\psi_{v}$ is tempered for any place $v$, and $\pi_{v}\in\Pi_{\psi_{v}}$ is tempered. This proves Part (1).
For Part (2), write $\psi=\boxplus_{i=1}^{m}\mu_{i}\otimes S_{b_{i}}$ and write
$\psi_{v}=\bigoplus_{i=1}^{m}\left(\bigoplus_{j\in I_{\mu_{i}}}\tau_{j}\otimes S_{a_{j}}\right)\otimes S_{b_{i}}$
for any finite place $v$ such that $\pi_{v}$ is unramified. Since $\pi_{v}\in\Pi_{\psi_{v}}$, Corollary 5.3 implies $\pi\in\Pi_{\phi_{\psi_{v}}}$ and $\phi_{\psi_{v}}$ is unramified. Therefore $a_{j}=1$ for any $j\in I_{\mu_{i}}$. On the other hand, assuming Conjecture 6.4, Lemma 6.5 implies the eigenvalues of $\tau_{j}(\mathrm{Frob}_{v})$ all have absolute value one. We conclude that
$\{{w_{v,1}},\ldots,{w_{N,1}}\}$ $=\sum_{i=1}^{m}\sum_{j\in I_{\mu_{i}}}{\dim(\tau_{j})}\left(\left\{\frac{b_{i}-1}{2},\frac{b_{i}-3}{2},\ldots,\frac{1-b_{i}}{2}\right\}\right)$
$=\sum_{i=1}^{m}\dim(\mu_{i})\left(\left\{\frac{b_{i}-1}{2},\frac{b_{i}-3}{2},\ldots,\frac{1-b_{i}}{2}\right\}\right),$
which is independent of the place $v$. Here $\dim(\tau_{j})$ and $\dim(\mu_{i})$ denotes the multiplicities. This completes the proof of Part (2) and the theorem. ∎
Based on the proof of Part (1) above and Theorem 6.2, we propose the following conjecture for general reductive groups $\mathrm{G}$.
Conjecture 6.12.
Let $\pi\in\mathcal{A}_{\mathrm{G},disc}$. Suppose there exists a finite place $v_{0}$ of $k$ such that $\pi_{v_{0}}$ is generic, then $\pi_{v}$ is generic for almost all places.
Note that in the case of $\mathrm{G}=\mathrm{G}_{n}$, this is proved in Theorem 6.2(2) as an application of enhanced Shahidi conjecture. We remark that in either Conjecture 6.12 or Theorem 6.2(2), we do not assume Conjecture 6.4. On the other hand, if we do assume Conjecture 6.4, then we obtain a stronger conclusion.
Theorem 6.13.
Assume Conjecture 6.4. Suppose that $\pi=\otimes\pi_{v}$ lies in a global Arthur packet $\Pi_{\psi}$ of $\mathrm{G}_{n}({\mathbb{A}}_{k})$ and that there exists a finite place $v_{0}$ such that $\pi_{v_{0}}$ is generic. Then $\pi_{v}$ is tempered for all places $v$.
Proof.
By Theorem 6.2(1), the global Arthur parameter is of the form $\psi=\boxplus_{i=1}^{m}\mu_{i}\otimes S_{1}$, and hence for any place $v,$ the local Arthur parameter $\psi_{v}$ is trivial on the Arthur-$\mathrm{SL}_{2}(\mathbb{C})$. On the other hand, assuming Conjecture 6.4, Lemma 6.5 implies that $\psi_{v}$ has bounded image when restricted to $W_{F}$. We conclude that for any place $v$, $\psi_{v}$ is tempered, and hence $\pi_{v}\in\Pi_{\psi_{v}}$ is also tempered. This completes the proof of the theorem. ∎
We remark that Theorem 6.13 improves the results in [Sha11, Theorem 6.1 and Conjecture 6.5], where $\pi$ is assumed to be locally generic (at every local place).
Finally, we recall [Clo07, Conjecture 5] which is also true under the assumption of Conjecture 6.4, as a corollary of Theorem 6.13.
Conjecture 6.14 ([Clo07, Conjecture 5]).
Let $\pi\in\mathcal{A}_{\mathrm{G},disc}$. Suppose there exists a finite place $v_{0}$ of $k$ such that $\pi_{v_{0}}$ is a Steinberg representation, then $\pi_{v}$ is tempered for all places $v$.
Theorem 6.15.
When $\mathrm{G}=\mathrm{G}_{n}$, Conjecture 6.4 implies Conjecture 6.14.
We remark that Clozel originally made Conjectures 6.6, 6.10 and 6.14 for $L^{2}$-automorphic representations, here for simplicity, we only considered the case of discrete spectrum. Note that by Langlands’ theory of Eisenstein series, the conjectures for $L^{2}$-automorphic representations can be reduced to those for the discrete spectrum.